### v2

# thanks: Full author list given at the end of the paper

Abstract
The study of the charmed baryons is crucial for investigating the strong and weak interactions in the Standard Model and for gaining insights into the internal structure of baryons. In an $e^{+}e^{-}$ experiment the lightest charmed baryon, $\Lambda_{c}^{+}$, can be produced in pairs through the single photon annihilation process. This process can be described by two complex electromagnetic form factors. The presence of a non-zero relative phase between these form factors gives rise to a transverse polarization of the charmed baryon and provides additional constraints on the dynamic parameters in the decays. In this article, we present the first observation of the transverse polarization of $\Lambda_{c}^{+}$ in the reaction $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, based on $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at center-of-mass energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The decay asymmetry parameters and strong phase shift in the decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, $\Sigma^{+}\pi^{0}$ are also simultaneously extracted from the joint angular distributions. These results are vital for understanding CPC\!P violation and its role in the matter-antimatter asymmetry of the Universe.
I Introduction
Although the most abundant constituent of matter is baryons, our understanding of them is less knowledgeable than that of mesons. The addition of one constituent quark in baryons compared to the quark-antiquark mesons introduces numerous challenges in theoretical calculations, making high-order and diquark correlation contributions difficult to predict accurately. One of the effective approaches to study the structure of baryons is to measure their electromagnetic form factors (EMFFs) [1]. In the $e^{+}e^{-}$ annihilation process, the time-like form factors are accessible, providing insights into the electromagnetic structure of the baryons [2]. For a spin-1/2 baryon ($B$), the process of $e^{+}e^{-}\to B\bar{B}$ is dominated by the single photon exchange. The cross-section can be described by the complex electric form factor $G_{E}$ and magnetic form factor $G_{M}$, which are functions of the four-momentum squared $q^{2}$ of the photon. The angular distribution of the produced baryon is characterized by two parameters: the angular parameter $\alpha_{0}$ and the phase difference between the EMFFs $\Delta\Phi$, where $\Delta\Phi$ is defined as $\text{arg}(G_{M})-\text{arg}(G_{E})$ [3]. The transverse polarization can be parameterized as
$P_{y}(\cos\theta_{0})=\frac{3}{2(3+\alpha_{0})}\sqrt{1-\alpha^{2}_{0}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi,$ (1)
where the factor $\frac{3}{2(3+\alpha_{0})}$ arises from normalization in spin density matrix of $\Lambda_{c}^{+}$. A non-vanishing $\Delta\Phi$ gives rise to the transverse polarization effect of the baryon, which depends on the momentum direction of the baryon ($\theta_{0}$) in the rest frame of $e^{+}e^{-}$. This polarization effect has been observed in the pair production of hyperons $\Lambda$ [4, 5, 6], $\Sigma^{0,+}$ [7, 8, 9] and $\Xi^{0,-}$ [10, 11, 12] by the BESIII Collaboration, which has led to the highest precision measurements of the decay parameters [13] of the hyperons. However, this effect has never been observed in charmed baryons.
The $\Lambda_{c}^{+}$ is the lightest charmed baryon, where a light up quark in the proton is replaced by a heavy charm quark. The study of the $\Lambda_{c}^{+}$ decay is of significant importance in investigating the fundamental theory of the strong and weak interactions. Similar to the pair production of strange hyperons, it is expected that the $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ production can also acquire polarization with respect to the normal of the production plane.
Figure 1: Generation and transmission of polarization in a weak decay $\Lambda_{c}^{+}\to BP$ [14].
In the weak decays of $\Lambda_{c}^{+}\to BP$, where $B$ and $P$ denote a spin-1/2 baryon and a pseudoscalar meson, respectively, the polarization of $\Lambda_{c}^{+}$, $\bm{\mathcal{P}_{\Lambda_{c}^{+}}}$, can be transferred to the subsequent decays and manifested in the angular distributions of the daughter baryons, as shown in Fig. 1. The polarization of the daughter baryon, denoted by $\bm{\mathcal{P}_{B}}$, is expressed as
$\small\bm{\mathcal{P}_{B}}=\frac{(\alpha+\bm{\mathcal{P}_{\Lambda_{c}^{+}}}\cdot\bm{\hat{n}})\bm{\hat{n}}+\beta(\bm{\mathcal{P}_{\Lambda_{c}^{+}}}\times\bm{\hat{n}})+\gamma\bm{\hat{n}}\times(\bm{\mathcal{P}_{\Lambda_{c}^{+}}}\times\bm{\hat{n}})}{1+\alpha\bm{\mathcal{P}_{B}}\cdot\bm{\hat{n}}}$, (2)
[…]
Various phenomenological models taking into account non-perturbative effects and non-factorizable contributions in quantum chromodynamics (QCD), have been developed to predict the decay asymmetries of the charmed baryons. The development of these models can be roughly divided into three stages. The first stage was in the last century, after CLEO Collaboration first measured $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ in 1990 [16]. Models such as covariant confined quark model (CCQM) [17, 18], the pole model [19, 20, 21, 22, 23], and current algebra (CA) [24, 25] were applied to predict decay asymmetry parameters. The decay asymmetry parameters $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ predicted by different models varied significantly, particularly with regard to uncertainty on the sign. The beginning of the second stage is marked by the BESIII Collaboration’s measurement of the decay asymmetry in $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays together with the first determination of $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ in 2019 [15]. The sign of $\alpha_{pK^{0}_{S}}$ is positive with a large uncertainty, while all predictions are negative. Another point is that the sign of $\alpha_{\Sigma^{0}\pi^{+}}$ has been confirmed. The Belle Collaboration results for $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ [26, 27] are in a good agreement with the flavor SU(3) symmetry. Subsequently, the BESIII Collaboration first reported the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay asymmetry parameters in 2024 [28], which is a pure $W$-exchange process. The non-perturbative nature and the problem of non-factorization in this decay channel make predictions very difficult. A persistent tension exists between the branching fraction and the decay asymmetry parameters, as theoretical predictions and experimental measurements do not align perfectly. To resolve this tension, the study of the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay introduces a valuable perspective as the strong phase shift, which describes the relative phase difference induced by strong interactions in the decay, has a crucial role and may significantly influence the decay asymmetry parameters. Following this, some theoretical works [29, 30, 31] have provided detailed discussions on this measurement and also introduced a strong phase shift, which significantly improves the description of the data. The most notable change due to the strong phase shift is seen for $\alpha_{pK^{0}_{S}}$, where this parameter changes from a negative value to one consistent with zero. This modification leads to a substantial change in the related calculations, making the precise measurement of $\alpha_{pK^{0}_{S}}$ crucial. In 2024, the LHCb Collaboration measured $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Lambda\pi^{+}}$ with high accuracy using the $\Lambda_{b}^{0}\to\Lambda_{c}^{+}\pi^{-}$ decay, which provides a large longitudinal polarization to $\Lambda_{c}^{+}$ in the decay process  [32]. These results agree with the latest theoretical predictions  [33]. It is clear that the value of $\alpha_{pK^{0}_{S}}$ should be negative rather than close to zero. The BESIII Collaboration can also independently measure $\alpha_{pK^{0}_{S}}$ using a large sample to confirm the sign of $\alpha_{pK^{0}_{S}}$. For $\alpha_{\Lambda\pi^{+}}$, the Belle Collaboration reported $\alpha_{\Lambda\pi^{+}}$ [26] with the same accuracy as the result from the LHCb Collaboration, but did not report $\beta_{\Lambda\pi^{+}}$ and $\gamma_{\Lambda\pi^{+}}$. All theoretical predictions and measurements about $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ are summarized in Table 1.
Table 1: Theoretical calculations and experimental measurements of the decay asymmetry parameters. The superscript $a$ means the model only considered SU(3) symmetry, and $b$ means considered the contribution of both SU(3) symmetry and SU(3) broken.
Predictions or experiments $\alpha_{pK^{0}_{S}}$ $\alpha_{\Lambda\pi^{+}}$ $\alpha_{\Sigma^{0}\pi^{+}}$ $\alpha_{\Sigma^{+}\pi^{0}}$
Körner(1992), CCQM [17] 0.10-0.10 0.70-0.70 $0.70$ $0.71$
Xu(1992), Pole [19] $0.51$ 0.67-0.67 $0.92$ $0.92$
Cheng, Tseng(1992), Pole [20] 0.49-0.49 0.96-0.96 $0.83$ $0.83$
Cheng, Tseng(1993), Pole [21] 0.49-0.49 0.95-0.95 $0.78$ $0.78$
Źencaykowski(1994), Pole [22] 0.66-0.66 0.99-0.99 $0.39$ $0.39$
Źencaykowski(1994), Pole [23] 0.90-0.90 0.86-0.86 0.76-0.76 0.76-0.76
Alakabha Datta(1995), CA [24] 0.91-0.91 0.94-0.94 0.47-0.47 0.47-0.47
Ivanov(1998), CCQM [18] 0.97-0.97 0.95-0.95 $0.43$ $0.43$
Sharma(1999), CA [25] 0.99-0.99 0.99-0.99 0.31-0.31 0.31-0.31
Geng(2019), SU(3) [34] $0.89^{+0.26}_{-0.11}$ $0.87\pm 0.10$ $0.35\pm 0.27$ $0.35\pm 0.27$
Zou(2020), CA [35] 0.75-0.75 0.93-0.93 0.76-0.76 0.76-0.76
Zhong (2022), SU(3)a [36] $0.51\pm 0.21$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Zhong (2022), SU(3)b [36] $0.29\pm 0.24$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Liu(2023), Pole [37] $0.81\pm 0.05$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Liu(2023), LP [37] $0.68\pm 0.01$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Geng(2023), SU(3) [29] $0.40\pm 0.49$ $0.75\pm 0.01$ $0.47\pm 0.02$ $0.47\pm 0.02$
Zhong(2024), TDA [31] $0.01\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), IRA [31] $0.03\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), TDA [33] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
Zhong(2024), IRA [33] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
CLEO(1990) [16] $1.0^{+0.4}_{-0.1}$
ARGUS(1992) [38] $0.96\pm 0.42$
CLEO(1995) [39] $0.94^{+0.21+0.12}_{-0.06-0.06}$ $0.45\pm 0.31\pm 0.06$
FOCUS(2006) [40] $0.78\pm 0.16\pm 0.19$
BESIII(2019) [15] $0.18\pm 0.43\pm 0.14$ $0.80\pm 0.11\pm 0.02$ $0.73\pm 0.17\pm 0.07$ $0.57\pm 0.10\pm 0.07$
BELLE(2022) [26, 27] $0.755\pm 0.005\pm 0.003$ $0.463\pm 0.016\pm 0.008$ $0.48\pm 0.02\pm 0.02$
LHCb(2024) [32] $0.754\pm 0.008\pm 0.006$ $0.785\pm 0.006\pm 0.003$
PDG Fit(2025) [41] $0.754\pm 0.010$ $0.768\pm 0.015$ $0.466\pm 0.018$ $0.484\pm 0.027$
The asymmetry in charge conjugation and parity symmetry (CPC\!P) lies at the heart of some of the most fundamental problems in physics. The presence of CPC\!P asymmetry has been established in meson decays [42, 43, 44]. Recently, the LHCb collaboration reported the first observation of direct CPC\!P asymmetry in a baryonic decay [45], $\Lambda_{b}^{0}\to pK^{-}\pi^{+}\pi^{-}$, marking a significant milestone in the study of CPC\!P violation. Meanwhile, ongoing efforts are exploring CPC\!P asymmetry in other baryonic processes [46, 47]. Despite these discoveries, the origin of CPC\!P asymmetry and a precise determination of its magnitude in specific processes remain open questions. Its size strictly depends on the strong and weak phase shift. In Cabbibo-favored (CF) decays of the $\Lambda_{c}^{+}$ baryon, there is no CPC\!P asymmetry, as the weak phase difference is expected to be zero. By relaxing the constraints of CPC\!P asymmetry , we can independently extract both the strong and weak phase shifts. The strong phase shift is an important input for theoretical calculations, helping to refine models of hadronic dynamics and improve our understanding of the underlying strong interaction. On the other hand, the weak phase provides a direct test of Standard Model (SM) predictions, offering valuable insights into potential CPC\!P asymmetry and verifying the consistency of weak interactions in charm decays.
In this paper, we observe for the first time the transverse polarization of the $\Lambda_{c}^{+}$ baryon in $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ and obtain the phase difference ($\Delta\Phi$) between the EMFFs at thirteen CM energies ranging from $4600$ to $4951$ MeV, with an overall significance exceeding $10.0\sigma$. The decay asymmetry parameters ($\alpha_{BP}$ and $\Delta_{BP}$, which then leads to $\beta_{BP}$ and $\gamma_{BP}$) and phase shifts between the $S$- and $P$-wave ($\delta_{p}-\delta_{s}$) are measured for four CF channels: $\Lambda_{c}^{+}\to pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$. The decay of $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, notable for its clean structure and large branching fraction, is included to improve the transverse polarization measurement. By measuring the processes in samples separated by baryon charge, the CPC\!P asymmetry parameters, strong phase shifts and weak phase shifts between different partial waves are also extracted in the CF processes. This allows precise testing of the SM and provides an essential input for the theoretical predictions. The analysis is performed based on a multidimensional angular analysis of the cascade decays of $\Lambda_{c}^{+}$, as described in Ref. [15]. The charge conjugation channels are included in the analysis and are implied in the context unless explicitly stated otherwise.
II Experimental apparatus and data sets
The Beijing Spectrometer III (BESIII) detector [48] records symmetric $e^{+}e^{-}$ collisions provided by the Beijing Electron Positron Collider (BEPCII) storage ring [49] in the CM energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [50, 51, 52]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps [53, 54, 55]. Data analyzed in this work were collected at thirteen energy points above the $\Lambda^{+}_{c}\bar{\Lambda}^{-}_{c}$ threshold. The integrated luminosity of the datasets analyzed in this work is estimated to be $6.4~\mathrm{fb}^{-1}$ [56, 57].
Simulated data samples produced with a geant4-based [58] Monte Carlo package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [59, 60]. The inclusive simulation includes the production of open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc [59, 60]. All particle decays are modelled with evtgen [61, 62] using branching fractions either taken from the Particle Data Group (PDG) [41], when available, or otherwise estimated with lundcharm [63, 64]. Final state radiation (FSR) from charged final state particles is incorporated using the photos package [65]. The phase space (PHSP) signal simulation is generated by simulating all allowed kinematic configurations within the entire phase space for the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, followed by the decay of $\Lambda_{c}^{+}$ into the signal final states and the inclusive decay of $\bar{\Lambda}_{c}^{-}$, which accounts for all possible decay modes. The signal simulation is based on the process model described in the helicity formalism with the decay asymmetry parameters measured in this work or cited from the PDG [41].
III Measurement Method
III.1 Event selection and $\Lambda_{c}^{+}$ signal in the beam-constrained mass spectrum
A single tag approach [66] is employed, where only one $\Lambda_{c}^{+}$ or $\bar{\Lambda}_{c}^{-}$ baryon is reconstructed in each event. Four two-body hadronic decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, $\Sigma^{+}\pi^{0}$, and one three-body decay, $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, are studied in this research following Refs. [66, 67]. The intermediate states, $K_{S}^{0}$, $\Lambda$, $\Sigma^{0}$, and $\Sigma^{+}$ are reconstructed by combining $\pi^{+}\pi^{-}$, $p\pi^{-}$, $\gamma\Lambda$, and $p\pi^{0}$ candidates, respectively.
Charged tracks detected in the MDC are required to satisfy $|\cos\theta|<0.93$, where $\theta$ is the polar angle defined relative to the symmetry axis of the MDC (the $z$-axis). Additionally, the tracks that do not come from a $K^{0}_{S}$ meson or $\Lambda$ baryon must have a distance of closest approach to the interaction point (IP) satisfying $|r_{z}|<10~\text{cm},\quad|r_{xy}|<1~\text{cm}$, where $r_{z}$ and $r_{xy}$ are the distances along the beam axis and in the transverse plane, respectively.
Particle identification (PID) for charged tracks combines measurements of the energy deposited in the MDC (dEE/dxx) and the flight time in the TOF to form likelihoods $\mathcal{L}(h)~(h=p,K,\pi)$ for each hadron $h$ hypothesis. Tracks are identified as protons if their likelihoods satisfy $\mathcal{L}(p)>\mathcal{L}(K)$ and $\mathcal{L}(p)>\mathcal{L}(\pi)$. Charged kaons and pions are identified by comparing the likelihoods of the respective hypotheses: $\mathcal{L}(K)>\mathcal{L}(\pi)$ and $\mathcal{L}(\pi)>\mathcal{L}(K)$, respectively.
Neutral showers are reconstructed in the EMC. Showers not associated with any charged track are identified as photon candidates. The deposited energy of each shower in the EMC must satisfy the following conditions: $E_{\text{barrel}}>25$ MeV for $|\cos\theta|<0.80$, $E_{\text{end-cap}}>50$ MeV for $0.86<|\cos\theta|<0.92$. The EMC time difference from the event start time is required to be less than 700 ns to suppress electronic noise and showers unrelated to the event. The $\pi^{0}$ candidates are reconstructed from photon pairs with invariant masses satisfying: $115~\text{MeV}/c^{2}<M(\gamma\gamma)<150~\text{MeV}/c^{2}$.
Candidates for $K^{0}_{S}$ and $\Lambda$ decays are reconstructed from $\pi^{+}\pi^{-}$ and $p\pi^{-}$ combinations, respectively. For these tracks, the distances of closest approach to the interaction point (IP) must satisfy: $|r_{z}|<20~\text{cm}$. There is no distance constraint required in the transverse plane. The charged pions in the $K^{0}_{S}$ reconstruction are not subject to the PID requirements described earlier, while the proton in the $\Lambda$ reconstruction is required to satisfy the PID criteria. The two final-state tracks are constrained to originate from a common decay vertex by requiring the vertex fit to satisfy: $\chi^{2}<100$. Furthermore, the decay vertex is required to be separated from the IP by a distance of at least twice the fitted vertex resolution. To select $K^{0}_{S}$ and $\Lambda$ candidates, the invariant masses of the combinations must satisfy: $487~\text{MeV}/c^{2}<M(\pi^{+}\pi^{-})<511~\text{MeV}/c^{2}$, $1111~\text{MeV}/c^{2}<M(p\pi^{-})<1121~\text{MeV}/c^{2}$. These mass windows correspond to approximately three times the standard deviation on either side of the known masses.
To suppress the peaking backgrounds for some decay channels, like $\Lambda_{c}^{+}\to\Lambda l^{+}\nu_{l}(l=e/\mu)$ background in the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, and $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, additional requirements are applied. Positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. For the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, it is required that $\mathcal{L}^{\prime}(e)<0.05$ and $\mathcal{L}(\pi)>0.001$ are satisfied to suppress the $\Lambda_{c}^{+}\to\Lambda e^{+}\nu_{e}$ background. For the $\Lambda_{c}^{+}\to\Lambda\mu^{+}\nu_{\mu}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, the information from the muon counter detector is utilized, and the penetrated depth of the $\pi^{+}$ candidates is required to be less than $5~\text{cm}$. A kinematic variable, the energy difference $\Delta E\equiv E_{\Lambda_{c}^{+}}-E_{\text{beam}}$, is used to suppress background, where $E_{\Lambda_{c}^{+}}$ is the total measured energy of the $\Lambda_{c}^{+}$ candidate, and $E_{\text{beam}}$ is the beam energy. For the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$, $pK^{0}_{S}$, $\Lambda\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays the requirements are $\Delta E\in(-29,26)$, $\Delta E\in(-21,18)$, $\Delta E\in(-23,21)$, and $\Delta E\in(-67,32)$, respectively. For the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, an additional photon can be selected from the process $\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and its $\Delta E$ is expected to be larger than zero. The requirement $\Delta E\in(-33,20)~\text{MeV}$ is imposed and can suppress this type of background efficiently.
A large background for the channel $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$ is from $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ combined with a random photon. This background, which may pose hidden dangers to the angular analysis, necessitates the application of a four-constraint kinematic fit to the $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$, where $\Sigma^{+}$ decays to proton and $\pi^{0}$, $\pi^{0}\to\gamma\gamma$. The quality of the fit, $\chi^{2}$, is calculated by constraining two $\pi^{0}$s, $\Sigma^{+}$, and $\Lambda_{c}^{+}$ to their world average masses quoted from PDG [41]. After optimization by $N^{2}_{S}/(N_{S}+N_{B})^{3/2}$ [68, 69], the requirement is set that the $\chi^{2}$ must be less than 12, where $N_{S}$ and $N_{B}$ represent the signal and background yields, respectively.
After applying the above selection criteria, the beam-constrained mass $M_{\text{BC}}\equiv\sqrt{E_{\text{beam}}^{2}/c^{4}-|\vec{p}_{\Lambda_{c}^{+}}|^{2}/c^{2}}$, is defined to identify signal candidates, and $\vec{p}_{\Lambda_{c}^{+}}$ is the momentum of the $\Lambda_{c}^{+}$ candidates calculated in the $e^{+}e^{-}$ rest frame. An unbinned maximum-likelihood fit is performed on the distribution of $M_{\text{BC}}$. The shapes of both correctly and mis-reconstructed signal events are modeled with the simulation, convolved with a Gaussian function to account for the resolution difference between the data and the simulation. The truth matching technique classifies events as correctly reconstructed if the angle between the reconstructed and true photon momentum directions from the simulation is below 10 degrees; otherwise, they are mis-reconstructed. The background shapes are described by an ARGUS function [70], except in the case of the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channel at CM energies larger than $4740$ MeV, where a second-order Chebyshev polynomial function is employed to model the background. The fit results are shown in Fig. 2 for all channels at $4600$ MeV. The obtained signal yields, combining all energy points, are $50083\pm 233$, $9619\pm 96$, $5742\pm 74$, $2487\pm 48$ and $1268\pm 37$ in the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$$pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$ channels, respectively, within the signal region of $M_{\text{BC}}\in[2.282,2.291]~\text{GeV}/c^{2}$.
Figure 2: Beam-constrained mass spectrum for $\Lambda_{c}^{+}$ signal at $4600$ MeV. Each panel shows a different signal channel as labeled. The Black points with error bars are data. The red, and black shaded region indicate the fitted signal and mis-reconstructed signal events. The yellow region represents the fitted combinatorial background.
III.2 Helicity system and angular distribution definitions
A helicity system and helicity angles are employed in constructing the amplitude for $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, $\Lambda_{c}^{+}\to BP$. Taking the channel $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ as an example, a kinematic diagram of this decay is shown in Fig. 3. For the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, the momenta of the $\Lambda_{c}^{+}$ and $\bar{\Lambda}_{c}^{-}$ baryons are defined in the CM system of the $e^{+}e^{-}$ pair, and $\theta_{0}$ is the polar angle of the $\Lambda_{c}^{+}$ in CM system. For the helicity system describing $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ decay, the momenta of the $\Lambda$ and the $\pi^{+}$ are defined in the $\Lambda_{c}^{+}$ frame. The angle $\phi_{1}$ is that between the $e^{+}\Lambda_{c}^{+}$ and $\Lambda\pi^{+}$ planes, and $\theta_{1}$ is the polar angle of the $\Lambda$ momentum defined in the rest frame of the $\Lambda_{c}^{+}$ baryon. For the helicity system describing the $\Lambda\to p\pi^{-}$, $\phi_{2}$ and $\theta_{2}$ adopt similar definitions. Then the formula of angular distributions related to this helicity frame based on the language of helicity amplitudes and Wigner-D function can be derived [15].
Figure 3: Definitions of the helicity frames and related angles for $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-},\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and $\Lambda\to p\pi^{-}$.
The joint angular distribution of the single-side decay chains can be derived from polarization transmission [3] or equivalently the helicity density matrix [71]. The angular distribution of the decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to\Lambda(\to p\pi^{-})\pi^{+})$ is shown below:
$\small\begin{split}&\frac{\text{d}\varGamma}{\text{d}\cos\theta_{0}~\text{d}\cos\theta_{1}~\text{d}\cos\theta_{2}~\text{d}\phi_{1}~\text{d}\phi_{2}}\\ &\propto 1+\alpha_{0}\cos^{2}\theta_{0}\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{p\pi^{-}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\sin\theta_{1}\sin\phi_{1}\cos\theta_{2}\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{p\pi^{-}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\cos\theta_{1}\sin\phi_{1}\sin\theta_{2}\\ &\times\sqrt{1-\alpha^{2}_{\Lambda\pi^{+}}}\cos(\Delta_{\Lambda\pi^{+}}+\phi_{2})\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{p\pi^{-}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\cos\phi_{1}\sin\theta_{2}\\ &\times\sqrt{1-\alpha^{2}_{\Lambda\pi^{+}}}\sin(\Delta_{\Lambda\pi^{+}}+\phi_{2})\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{\Lambda\pi^{+}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\sin\theta_{1}\sin\phi_{1}\\ &+\alpha_{0}\alpha_{p\pi^{-}}\alpha_{\Lambda\pi^{+}}\cos^{2}\theta_{0}\cos\theta_{2}\\ &+\alpha_{p\pi^{-}}\alpha_{\Lambda\pi^{+}}\cos\theta_{2},\end{split}$ (3)
where the $\alpha_{p\pi^{-}}$ denotes the decay asymmetry parameter in the weak hadronic decay $\Lambda\to p\pi^{-}$. The parameter $\Delta_{\Lambda\pi^{+}}$ is the phase shift between the helicity amplitudes of $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ satisfying $\beta_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\sin\Delta_{\Lambda\pi^{+}},\gamma_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\cos\Delta_{\Lambda\pi^{+}}$. Formulas of other two-body decays can be found in Ref. [15].
Although the analysis of the angular distribution of these four two-body CF channels provides effective constraints, the three-body channel $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, with a low background and large production rate, brings beneficial additional sensitivity. Benefiting from the transition amplitude analysis reported by the LHCb Collaboration [72], the angular distribution of decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to pK^{-}\pi^{+})$ can be expressed explicitly as:
$\small\begin{split}&\frac{\text{d}\mathcal{\varGamma}}{\text{d}\cos\theta_{0}~\text{d}\cos{\theta_{1}}~\text{d}\phi_{1}~\text{d}\phi_{2}}\\ &\propto I_{0}(\kappa)\\ &+I_{0}(\kappa)\cos^{2}\theta_{0}\\ &+I_{0}(\kappa)\sqrt{1-\alpha^{2}_{0}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\sum_{j}R_{2j}(\phi_{1},\theta_{1},\phi_{2})\alpha_{j}(\kappa),\end{split}$ (4)
where $\phi_{1},\theta_{1},\phi_{2}$ are the helicity angles for the intermediate state ($K^{-}\pi^{+}$), $R_{ij}(\phi_{1},\theta_{1},\phi_{2})$ represents the 3-dimension $Z$-$Y$-$Z$ rotation matrix and $\kappa$ denotes a set of Dalitz-plot variables $(M(pK^{-})^{2},M(K^{-}\pi^{+})^{2})$. The $\vec{\alpha}(\kappa)$ and $I_{0}(\kappa)$ are model-agnostic representations for polarization dependence of the decay rate and the total differential decay rate over $\kappa$, respectively, which are taken from the $\Lambda_{c}^{+}$ polarimetry with the amplitude model both measured by LHCb Collaboration [72, 73].
III.3 Fit of the joint angular distributions
In this analysis, the free parameters of $\alpha_{0}$, $\Delta\Phi$, $\alpha_{BP}$, and $\Delta_{BP}$, which describe the angular distributions for each CM energy are determined through a simultaneous unbinned maximum log-likelihood fit. The $\alpha_{0}$ and $\Delta\Phi$ are specific to the invididual energy point. The likelihood function is constructed from the joint probability density function (PDF) by
$\mathcal{L}_{\text{total}}=\sum^{\text{energy}}\mathcal{L}_{\text{data}}=\prod_{i=1}^{N_{\text{data}}}f_{s}(\vec{\xi}_{i}).$ (5)
Here, $f_{s}(\vec{\xi}_{i})$ is the PDF of the signal process, $N_{\text{data}}$ is the number of data events and $i$ is the event index. The signal PDF $f_{s}(\vec{\xi}_{i})$ is formulated as
$f_{s}(\vec{\xi}_{i})=\frac{\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}}{\int\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}\text{d}\vec{\xi}_{i}}~,$ (6)
where $\vec{\xi}_{i}$ denotes the kinematic observables, including the helicity angles and Daltiz-plot variables, while $\vec{\eta}$ denotes the free parameters to be determined. The $M(\vec{\xi}_{i};\vec{\eta})$ is the total decay amplitude, and $\epsilon(\vec{\xi}_{i})$ is the detection efficiency, parameterized in terms of the kinematic variables $\vec{\xi}_{i}$. It is important to note that all $\vec{\xi}_{i}$ are calculated using four-momentum updated by kinematic fit, which constrains the invariant mass of all resonances in the signal decay to their world average masses [41]. The background contribution to the joint likelihood is subtracted based on the calculated likelihoods using data in $M_{\text{BC}}\in[2.250,2.270]~\text{GeV}/c^{2}$. For the mis-reconstructed signal events, the contribution to the likelihood is based on the signal MC simulation. The integration of the normalization factor is calculated with a large PHSP signal simulation samples as $\int\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}\text{d}\vec{\xi}_{i}=\frac{1}{N_{\text{gen}}}\sum_{k_{\text{sim}}}^{N_{\text{sim}}}|M(\vec{\xi}_{k_{\text{sim}}};\vec{\eta})|^{2}$, where $N_{\text{gen}}$ is the total number of the generated events, $N_{\text{sim}}$ is the number of the events surviving all selection criteria and $k_{\text{sim}}$ is the event index.
The negative log-likelihood with background subtraction over the five signal channels and data samples collected from thirteen energy points is minimized using the Minuit [74] algorithms. The decay asymmetry parameters of the $\Lambda$, $\Sigma^{+}$, and $\Sigma^{0}$ baryons are fixed according to the PDG [41]. In the nominal fit, the parameters $\alpha_{BP}$ and $\Delta_{BP}$ are set to $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$, respectively, in accordance with the assumption of CPC\!P conservation. Here, $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ represent the parameters associated with the antiparticle $\bar{\Lambda}_{c}^{-}$ decays. The results of the fit are presented in Fig. 4. An alternative test is also performed where the assumption of CPC\!P conservation are relaxed and $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ are treated as free parameters in the fit.
Figure 4: Fit results of helicity angles combining all energy points. They are angular distribution about $pK_{S}^{0}$, $pK^{-}\pi^{+}$, $\Lambda\pi^{+}$, $\Sigma^{+}\pi^{0}$, and $\Sigma^{0}\pi^{+}$, respectively. Black points with error bars are data; the yellow shaded region indicates combinatorial background events and black shaded region is the mis-reconstructed signal events.
III.4 Systematic uncertainties
The systematic uncertainties originate from the differences between simulations and real physical processes leading to deviations in detection efficiencies. In this part, the absolute systematic uncertainties are estimated. The numerical values are summarized in Table 2, and the individual sources are discussed in the rest of this section.
  • Selection criteria. The systematic uncertainties regarding selection criteria primarily arise from the reconstruction of final states, $\Delta E$ requirement, and $M_{\text{BC}}$ signal selection. The reconstruction efficiency of proton, pion, and kaon tracks is studied with control samples of $^{+}e^{-}\to p\bar{p}\pi^{+}\pi^{-}, e+e$$^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, and $J/\psi\to K^{0}_{S}K^{\pm}\pi^{\mp}$ decays, respectively. The $\pi^{0}$ meson reconstruction efficiency is studied with the $\psi(3686)\to\pi^{0}\pi^{0}J/\psi$ and $e^{+}e^{-}\to\omega\pi^{0}$ decays. The control channels for the $K^{0}_{S}$ reconstruction efficiency are the $J/\psi\to\phi K^{0}_{S}K^{\mp}\pi^{\pm}$ and $J/\psi\to K^{*}(892)^{\pm}K^{\mp}$ decays. For the reconstruction efficiency of the $\Lambda$ baryon the $J/\psi\to\bar{p}K^{+}\Lambda$ and $J/\psi\to\Lambda\bar{\Lambda}$ decay samples are used [75]. The difference in reconstruction efficiency for photons between data and MC simulation is studied with the $J/\psi\to\pi^{+}\pi^{-}\pi^{0}$ decay and found to be $0.5\%$ and $1.5\%$ in the barrel and endcap of the EMC. A randomly generated efficiency correction factor is used to adjust the PHSP simulation. Requirements relating to the MUC depth, probability of electron, and pion assumption are studied with a control sample of $\Lambda^{+}_{c}\to\Lambda\pi^{+}$ decays. The potential differences in event selection efficiencies for data and MC simulation are taken into account in the fit described in the previous section via reweighting the PHSP sample. The uncertainties related to the $\Delta E$ and $M_{\text{BC}}$ requirements are evaluated by smearing the signal simulation with a Gaussian resolution function. The changes in the fit results based on new survived signal simulation are taken to be systematic uncertainties.
  • Performance of truth matching. The maximum difference between the matching condition with 9 or 11 degree and the nominal result is considered as systematic uncertainty.
  • Signal components. The amplitude-model related uncertainties are computed using the alternative models in Ref. [73]. With alternative polarimeter field inputs, the angular fit is repeated and the standard deviations and difference of mean value are assigned to the model uncertainties. In Ref. [72], the statistical uncertainties are estimated by varying the parameter values in the default model. These parameters are sampled from a Gaussian function, whose $\mu$ and $\sigma$ are given by the central and error values from the default model fit results in Ref. [73]. The polarimeter vector fields are reproduced for each varied parameter set. The angular fit procedure is repeated using the corresponding inputs, and the statistical component is given by taking the standard deviations and the difference of the mean value.
  • Background components. The background shape is extracted from the data events in the $M_{\text{BC}}$ sideband regions. To estimate the possible difference between the $M_{\text{BC}}$ sideband and signal regions, a weight factor is derived by comparing the $M(K^{-}\pi^{+})^{2}$ distributions in the $M_{\text{BC}}$ sideband and signal regions of inclusive simulation. Data events in the $M_{\text{BC}}$ sideband region are reweighted and the fit procedure is repeated. The difference is then taken as the uncertainty from background model. Background fractions are fixed according to the fit results from the nominal fit. To estimate systematic uncertainties, the background fraction for each energy point is varied by $\pm 1\sigma$ and the fit procedure is repeated. The quadratic sum of the largest variation from each fraction are assigned as systematic uncertainties from the background fraction. The mis-reconstructed model is examined using an alternative signal simulation produced with new input parameters $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ which are changed within $\pm 1\sigma$ of the results presented here. The uncertainty due to the combinational background model is estimated by varying the relative weights between $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs and other hadronic events based on the uncertainties of their cross-section ratio. The relevant systematic uncertainties from background size are examined by repeating the fits with an alternative background size obtained from the Gaussian sampling of the fitted parameters. The series of fitted parameters, assumed to follow a Gaussian distribution, are used to estimate the systematic uncertainty for the background size. The uncertainty is determined by summing the fitted Gaussian resolution and the deviation of the fitted Gaussian mean from the nominal value.
  • Fixed parameters. The systematic uncertainty due to the input parameters is evaluated by varying these parameters using a Gaussian sampling method. For each parameter, the obtained results are expected to follow a Gaussian distribution. The sum of the fitted Gaussian resolution and the difference between the fitted Gaussian mean and the nominal result is taken as the systematic uncertainty.
Table 2: Absolute systematic uncertainties in transverse polarization and decay asymmetry parameters (in percent), where SEL, TMP, SIG, BKG, and PAR in the table represents the sources from Selection criteria, Performance of truth matching, Signal components, Background components, and Fixed parameters, respectively. Some negligible terms are denoted by “”. The $\alpha_{0}$ and $\Delta\Phi$ are given as ranges because they vary for different energy points. The total uncertainty is calculated as the sum in quadrature of all components.
Parameters SEL TMP SIG BKG PAR Total
$\alpha_{0}$ $\leq 0.2$ $\leq 0.1$ $\leq 0.1$ 0.11.90.1-1.9 $\leq 0.1$ 0.11.90.1-1.9
$\Delta\Phi$ $\leq 0.3$ $\leq 0.1$ 0.45.80.4-5.8 0.62.50.6-2.5 $\leq 0.6$ 0.95.80.9-5.8
$\alpha_{pK^{0}_{S}}$ $0.1$ $1.1$ $2.9$ $0.1$ $3.1$
$\alpha_{\Lambda\pi^{+}}$ $0.8$ $0.3$ $0.3$ $0.9$
$\Delta_{\Lambda\pi^{+}}$ $0.9$ $0.8$ $0.8$ $0.1$ $1.4$
$\alpha_{\Sigma^{0}\pi^{+}}$ $0.3$ $0.1$ $0.8$ $0.2$ $0.9$
$\Delta_{\Sigma^{0}\pi^{+}}$ $1.5$ $1.0$ $2.2$ $0.3$ $2.9$
$\alpha_{\Sigma^{+}\pi^{0}}$ $1.1$ $1.2$ $1.4$ $2.2$
$\Delta_{\Sigma^{+}\pi^{0}}$ $1.2$ $1.2$ $3.7$ $0.3$ $4.0$
IV Results
IV.1 Transverse polarization
The results of the angular distribution fit, the $\alpha_{0}$, the phase difference $\Delta\Phi$, its significance, as well as the largest transverse polarization $\mathcal{P}_{y}^{max}$, are summarized in Table 3. The EMFF-related $\alpha_{0}$ is consistent with previous results but more precise. The ratio of $|G_{E}/G_{M}|$ is extracted using the equation $|G_{E}/G_{M}|^{2}(1-v^{2})=(1+\alpha_{0})/(1-\alpha_{0})$, where $v$ is the velocity of the $\Lambda^{+}_{c}$ in CM system, normalized to the speed of light. The ratios are compared with the previous BESIII measurements and theoretical predictions, as shown in Fig. 5. The measurements of $\sin\Delta\Phi$ are also presented in Fig. 5, with the theoretical predictions [76, 77] and other experimental results [15, 78]. There is some inconsistency between the measurements and the theoretical predictions.
Table 3: Results for $\alpha_{0}$, $\Delta\Phi$, and transverse polarization of $\Lambda_{c}^{+}$ at 13 energy points, where the first uncertainties are statistical and the second are systematic. In the table, SLSL means the significance level of non-zero $\Delta\Phi$, and $\mathcal{P}_{y}^{max}$ is the largest transverse polarization.
$\sqrt{s}~[\text{MeV}]$ $\alpha_{0}$~in previous work [15, 78] $\alpha_{0}$~in this work $\Delta\Phi~[\text{rad}]$ SLSL $\mathcal{P}_{y}^{max}$
$4600$ $0.20\pm 0.04\pm 0.02$ $0.226\pm 0.030\pm 0.004$ $0.100\pm 0.069\pm 0.009$ $2.2~\sigma$ $0.026\pm 0.018\pm 0.002$
$4612$ $0.26\pm 0.09\pm 0.01$ $0.160\pm 0.083\pm 0.004$ $0.146\pm 0.162\pm 0.030$ $1.1~\sigma$ $0.038\pm 0.042\pm 0.008$
$4628$ $0.21\pm 0.04\pm 0.01$ $0.181\pm 0.038\pm 0.001$ $0.371\pm 0.082\pm 0.012$ $6.8~\sigma$ $0.095\pm 0.020\pm 0.003$
$4641$ $09\pm 0.05\pm 0.01 0.0$$060\pm 0.039\pm 0.003$ $0.398\pm 0.073\pm 0.015$ $7.6~\sigma$ $0.099\pm 0.017\pm 0.004$
$4661$ $02\pm 0.05\pm 0.01 0.0$$008\pm 0.044\pm 0.003$ $0.496\pm 0.088\pm 0.021$ $8.5~\sigma$ $0.119\pm 0.019\pm 0.005$
$4682$ $0.15\pm 0.03\pm 0.01$ $0.102\pm 0.029\pm 0.003$ $0.502\pm 0.054\pm 0.021$ $14.1~\sigma$ $0.116\pm 0.011\pm 0.005$
$4699$ $0.34\pm 0.07\pm 0.01$ $0.305\pm 0.055\pm 0.010$ $0.545\pm 0.114\pm 0.028$ $7.1~\sigma$ $0.112\pm 0.021\pm 0.007$
$4740$ $0.49\pm 0.16\pm 0.03$ $358\pm 0.126\pm 0.008 0.0$$097\pm 0.190\pm 0.016$ $4~\sigma 0.0$$020\pm 0.039\pm 0.004$
$4750$ $0.42\pm 0.10\pm 0.01$ $0.347\pm 0.079\pm 0.004$ $0.316\pm 0.142\pm 0.019$ $3.1~\sigma$ $0.065\pm 0.029\pm 0.005$
$4781$ $0.17\pm 0.07\pm 0.01$ $0.157\pm 0.062\pm 0.007$ $0.395\pm 0.126\pm 0.028$ $5.1~\sigma$ $0.090\pm 0.027\pm 0.007$
$4843$ $0.38\pm 0.10\pm 0.01$ $0.282\pm 0.089\pm 0.019$ $0.385\pm 0.153\pm 0.034$ $3.6~\sigma$ $0.082\pm 0.031\pm 0.008$
$4918$ $0.62\pm 0.17\pm 0.01$ $0.612\pm 0.150\pm 0.019$ $0.423\pm 0.272\pm 0.024$ $1.9~\sigma$ $0.067\pm 0.043\pm 0.013$
$4951$ $0.63\pm 0.21\pm 0.01$ $0.744\pm 0.179\pm 0.007$ $0.700\pm 0.392\pm 0.058$ $1.8~\sigma$ $0.086\pm 0.050\pm 0.030$
Figure 5: Comparison of $|G_{E}/G_{M}|$ and $\sin\Delta\Phi$ between prediction and this work. Red points with error bars correspond to this work, and the gray ones are previous measurement results from BESIII [15, 78]. The purple dashed line with sky blue shadow is predicted from Ref. [76], and the blue solid line represents the predicted results with Ref. [77].
Of particular interest is the fact that the $\Lambda_{c}^{+}$ baryon is significantly polarized with respect to the normal of the production plane. The size of polarization is strongly dependent on $\Lambda_{c}^{+}$ momentum direction $\cos\theta_{0}$ with a maximum value of approximately 12%. To illustrate the polarization effect, the moment $\textless\sin\theta_{1}\sin\phi_{1}\textgreater$ is calculated as the average of $\sin\theta_{1}\sin\phi_{1}$ within a certain range of $\theta_{0}$. It is expected to be proportional to $\alpha_{BP}P_{y}(\cos\theta_{0})/(1+\alpha_{0}\cos^{2}\theta_{0})$. We divide the range of $\cos\theta_{0}$, (1,1)(-1,1), into eight bins, and calculate the moment for both data and reweighted simulation at energy points where the significance of $\Delta\Phi$ is larger than $5\sigma$, as shown in Fig. 6. To compute the significance, we perform a floating fit for $\Delta\Phi$ and a fixed fit for $\Delta\Phi=0$ to obtain the likelihood values and the degrees of freedom, from which we derive the significance of non-zero $\Delta\Phi$. The sign($\alpha_{BP}$) is used to avoid cancellation between charge-conjugated processes. The significance is reported in Table 3.
Figure 6: The moment $\textless\text{sign}(\alpha_{BP})\sin\theta_{1}\sin\phi_{1}\textgreater$ distribution to show the effect of tranverse polarization. It depends on $\cos\theta_{0}$ for $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channels and is drawn using points with significance of $\Delta\Phi$ larger than $5\sigma$. Points with error bars correspond to data and red band represents the simulation reweighted according to the fit results with total uncertainties.
IV.2 Decay asymmetry and SM test
As the SM predicts no CPC\!P asymmetry in these decay channels, we assume CPC\!P symmetry, meaning the decay asymmetry parameters $\alpha_{BP}=-\bar{\alpha}_{BP}$ and $\Delta_{BP}=-\bar{\Delta}_{BP}$ in the angular distribution analysis. In addition to the transverse polarization, the decay asymmetry parameters are obtained and shown in Fig. 7. In the SM, the amplitude for a spin-$1/2$ baryon decaying into a spin-$1/2$ baryon and a spin-0 meson can be written as $\mathcal{M}=i\bar{u}_{f}(A-B\gamma_{5})u_{i}$, where $A$ and $B$ are constants, that reflect the $S$ and $P$-wave amplitudes with different orbital angular momentum in weak decays, and directly reflect the mechanism of different weak decays. The $u_{i}$ and $\bar{u}_{f}$ are spinors describing the initial and final baryons [19]. The calculated method of $|A|$ and $|B|$ has been applied in $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ [28], successfully promoting the development of effective models [30, 31, 33]. In this work, $A$ and $B$ can also be derived, and their two-dimensional displays are shown in Fig. 8. We provide only the density plots and do not present numerical results due to the nonlinear issues encountered during the calculation, which makes it difficult to provide reliable uncertainty estimates.
Figure 7: The comparison about predictions and measurements of decay asymmetry parameters. Gray points are theoretical predictions [17, 19, 20, 21, 23, 22, 24, 18, 25, 34, 35, 36, 37, 29, 30, 31, 33] and experimental measurements [16, 38, 39, 40, 15, 26, 27, 32, 41] are marked in blue. Red points and yellow shadows represent the results obtained in this work. In the legend, the names of the theoretical studies and experimental measurements match the colors used in the figures, arranged in chronological order. For each theoretical study, the main theoretical models used are also specified.
Figure 8: The two-dimensional density of $|A|$ and $|B|$. Markers are theoretical predictions [19, 20, 21, 18, 35, 37, 34, 36, 31, 33] and experimental measurements [39], and shadows representing the distribution of $|A|$ and $|B|$ obtained in this work, $G_{F}$ is Fermi coupling constant, $c$ and $d$ means whether the model uses the results from Belle experiment with the BFs and decay asymmetry parameters of $\Xi_{c}^{0}\to\Xi^{0}\pi^{0}$, $\Xi^{0}\eta$, and $\Xi^{0}\eta^{\prime}$. The contours correspond to 68.2%, 95.4%, 99.7% confidence levels, respectively.
A secondary fit is also performed where the CPC\!P constraint is removed and is used to test the SM. The $\alpha$-induced CPC\!P observable, weak and strong phase shifts are constructed in $A^{\alpha_{BP}}_{C\!P}=\frac{\alpha_{BP}+\bar{\alpha}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\phi_{C\!P}=\frac{\beta_{BP}+\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\Delta_{S}=\frac{\beta_{BP}-\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, can be extracted for $\Lambda_{c}^{+}$ decay channels, where $\beta_{BP}=\sqrt{1-\alpha_{BP}^{2}}\sin\Delta_{BP}$, $\phi_{C\!P}$ and $\Delta_{S}$ denote the weak interaction phase shift between $P$-wave and $S$-wave introduced by phase angle in the Cabibbo-Kobayashi-Maskawa (CKM) matrix and the strong phase shift. The results are summarized in Table 4.
Table 4: Results of decay observables, where $\langle\alpha_{BP}\rangle$, $\langle\Delta_{BP}\rangle$, and $\delta_{p}-\delta_{s}$ are measured based on the assumption of CPC\!P conservation, while others are under the assumption of CPC\!P violation. Some negligible terms are denoted by “”. All angular values in the table are expressed in radians.
Parameter $\Lambda_{c}^{+}\to pK_{S}^{0}$ $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$
$\langle\alpha_{BP}\rangle$ $0.918^{+0.133}_{-0.082}\pm 0.031$ $0.790\pm 0.032\pm 0.009$ $0.502\pm 0.080\pm 0.009$ $0.590\pm 0.049\pm 0.022$
$\langle\Delta_{BP}\rangle$ $0.637\pm 0.444\pm 0.014$ $2.190\pm 0.730\pm 0.029$ $1.901\pm 0.603\pm 0.040$
$\langle\beta_{BP}\rangle$ $0.365^{+0.173}_{-0.246}\pm 0.010$ $0.704^{+0.143}_{-0.480}\pm 0.015$ $0.764^{+0.051}_{-0.237}\pm 0.018$
$\langle\gamma_{BP}\rangle$ $0.637^{+0.103}_{-0.202}\pm 0.011$ $0.502^{+0.591}_{-0.303}\pm 0.021$ $0.262^{+0.478}_{-0.383}\pm 0.031$
$\delta_{p}-\delta_{s}$ $2.71^{+0.28}_{-0.17}\pm 0.02$ $19^{+0.49}_{-0.13}\pm 0.02 2.2$$23^{+0.19}_{-0.06}\pm 0.03$
$A^{\alpha_{BP}}_{C\!P}$ $079^{+0.115}_{-0.101}\pm 0.019 0.0$$002\pm 0.047\pm 0.017$ $206^{+0.188}_{-0.156}\pm 0.028 0.0$$086\pm 0.081\pm 0.085$
$\text{tan}\phi_{C\!P}$ $0.232\pm 0.242\pm 0.025$ $393\pm 0.651\pm 0.042 0.0$$007\pm 0.474\pm 0.034$
$\text{tan}\Delta_{s}$ $0.475\pm 0.242\pm 0.029$ $1.411\pm 0.672\pm 0.062$ $1.297\pm 0.478\pm 0.068$
V Discussion
The effect of transverse polarization of $\Lambda_{c}^{+}$ baryons is discovered for the first time by analyzing $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at the CM energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The non-zero transverse polarization indicates the phase shift of the non-zero EMFF. For the first time, we scanned the phase shifts at multiple energy points and demonstrated their differences from theoretical predictions. In 2021, Ref. [76] used modified vector meson dominance (VMD) model considering vector charmed mesons and their excitations $J/\psi$, $\psi(2S)$, $\psi(3770)$, $\psi(4040)$, $\psi(4160)$, and $\psi(4415)$. They then constructed and fitted the Born cross section of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, since it was studied in 2021, where the data they fitted comes from the Belle Collaboration [79] and the BESIII Collaboration [66]. As the Belle Collaboration provided the main lineshape, the fitting is dominated by Belle’s results. The fitting results well describe the Born cross section of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$. However, Belle’s results showed a significant enhancement of $Y(4630)$, which was refuted by BESIII’s more precise results fifteen years later [78]. Based on this new result, another work [77] also based on the VMD model using resonances, $\psi(4500)$, $\psi(4660)$, $\psi(4790)$, and $\psi(4900)$, describes the Born cross section very well. Due to higher precision and the absence of abnormal $Y(4630)$ contribution, the fitted results describe the oscillation behavior of $|G_{E}/G_{M}|$ with energy points measured by BESIII Collaboration. Unfortunately, the results of these two works [76, 77] cannot describe the phase shifts between $G_{E}$ and $G_{M}$ very well. The theoretical predictions still show a significant deviation from the experimental measurements. Therefore, the stucture of $\Lambda_{c}^{+}$ and the mechanism of its production in the $e^{+}e^{-}$ collider remains a mystery.
The transverse polarization of the $\Lambda_{c}^{+}$ baryon provides an opportunity to measure the decay asymmetry parameters of the $\Lambda_{c}^{+}$ decay. The golden two-body decays $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ have been studied. The parameter $\alpha_{pK^{0}_{S}}$ has changed its sign, replacing the previous results [15]. The value of $\alpha_{pK^{0}_{S}}$ is now consistent with most theoretical calculations. The decay asymmetry parameters $\alpha_{\Sigma^{+}\pi^{0}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ support the prediction under $SU(3)$ flavor symmetry but $\alpha_{\Sigma^{+}\pi^{0}}$ is more than $2\sigma$ different than the measurements from Belle [27]. Due to the cascading decay in $\Lambda_{c}^{+}\to\Lambda\pi^{+},~\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$, we also obtain the value of $\Delta_{BP}$, and use this to derive $\beta_{BP}$ and $\gamma_{BP}$. This also enables us to extract the $S$- and $P$-wave phase shift, $\delta_{p}-\delta_{s}$. The results of phase shifts in this work are consistent with recent predictions [30, 31, 33] following the derived method from the LHCb Collaboration [32], and we use the convention that $sign=+1$ as in Ref. [14]. We have also derived the dynamic parameters $|A|$ and $|B|$ for weak interactions. Previous theoretical studies on the decay of charmed baryons have also predicted $|A|$ and $|B|$, but most of them did not account for the contribution of phase shifts to the decay asymmetry parameters, which complicates the accurate prediction of these parameters. Only a few recent works have emphasized the impact of phase shifts and predicted $|A|$ and $|B|$ [30, 31, 33]. Clearly, the reliability of their predictions is improving. The results presented here supersede those in Refs [15, 78].
A strong phase shift can be significant in physical processes and theoretically can greatly improve the CPC\!P asymmetry prediction [80, 29, 31]. In our work, we calculated the $\delta_{p}-\delta_{s}$ in three channels and demonstrated the existence of a large phase shift in charmed baryon decays. Such a large phase shift can be induced by re-scattering processes and loop effects, and further consideration of the role of final state interactions is warranted [29]. The decays measured in this work are all CF processes, where weak phase shifts are zero. Although the CPC\!P asymmetry phenomenon is impossible in the CF processes, it can still be transferred to other Cabibbo-suppressed processes through symmetry in physics. As discussed in Ref. [80], the re-scattering effects can enhance CPC\!P asymmetry by an order of magnitude, which makes it promising to observe CPC\!P asymmetry effect in baryon decays. In addition, based on the separate measurements of the positive and negative processes, we also test $A^{\alpha_{BP}}_{C\!P}$, strong and weak phase shifts. They are related by $A^{\alpha_{BP}}_{C\!P}=-\tan(\delta_{P}-\delta_{S})\tan(\xi_{P}-\xi_{S})$, and the results show that $A^{\alpha_{BP}}_{C\!P}$ and weak phase shifts are consistent with zero, which aligns with the expectations of the SM.
In conclusion, the observation of the transverse polarization effect of the $\Lambda_{c}^{+}$ with the BESIII detector represents a complete description in charmed baryon production in $e^{+}e^{-}$ annihilations, contributing to our understanding of its production properties. Our analysis of decay asymmetry parameters for $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ not only updates and validates previous measurements but also reveals significant strong phase shifts, providing valuable constraints on various theoretical models. Furthermore, the systematic study of production and decay properties provides a comprehensive dataset that refines global fits in flavor physics, enhances the understanding of CPC\!P violation in the charm sector, and aligns with current SM predictions. These findings, combined with the precise measurement of polarization dynamics, open new avenues for investigating hadronic interactions, spin effects, and charm quark dynamics. This work establishes a robust framework for future experimental programs at Belle II, LHCb, STCF and other facilities, setting the stage for deeper insights into the non-perturbative aspects of QCD and the exploration of potential new physics.
Data availability
The data analyzed in this study are not publicly available due to collaboration policies. Access to the underlying datasets may be granted upon reasonable request and with permission from the BESIII Collaboration by contacting besiii-publications@ihep.ac.cn.
Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819, 12422504, 12365015; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; Fundamental Research Funds for the Central Universities, Lanzhou University under Contracts Nos. lzujbky-2025-ytB01, lzujbky-2023-stlt01, lzujbky-2023-it32, University of Chinese Academy of Sciences; The Natural Science Foundation of Inner Mongolia Autonomous Region No. 2023QN01011; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; ERC under Contract No. 758462; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; STFC (United Kingdom); Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.