### v3

# thanks: Full author list given at the end of the paper

Abstract
The study of the charmed baryons is crucial for investigating the strong and weak interactions in the Standard Model and for gaining insights into the internal structure of baryons. In an $e^{+}e^{-}$ experiment the lightest charmed baryon, $\Lambda_{c}^{+}$, can be produced in pairs through the single photon annihilation process. This process can be described by two complex electromagnetic form factors. The presence of a non-zero relative phase between these form factors gives rise to a transverse polarization of the charmed baryon and provides additional constraints on the dynamic parameters in the decays. In this article, we present the first observation of the transverse polarization of $\Lambda_{c}^{+}$ in the reaction $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, based on $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at center-of-mass energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The decay asymmetry parameters in the decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ are simultaneously extracted from the joint angular distributions. From these parameters, both the weak and strong phase shifts are extracted, and several CPC\!P observables are tested. The obtained results show that the transverse polarization of the produced baryon significantly enhances the capability of searching for CPC\!P-violating effects, opening a novel way for exploring CPC\!P violation in the charmed baryon sector.
I Introduction
Although baryons populate with a significant abundance in the physical world surrounding us, our understanding of them is less knowledgeable than that of mesons. The addition of one constituent quark in baryons compared to the quark-antiquark mesons introduces numerous challenges in theoretical calculations, making high-order and diquark correlation contributions difficult to predict accurately. One of the effective approaches to study the structure of baryons is to measure their electromagnetic form factors (EMFFs) [Schonning:2023mge, Huang:2021xte, Pacetti:2014jai]. In the $e^{+}e^{-}$ annihilation process, the time-like form factors are accessible, providing insights into the electromagnetic structure of the baryons [Huang:2021xte]. For a spin-1/2 baryon ($B$), the process of $e^{+}e^{-}\to B\bar{B}$ is dominated by the single photon exchange. The cross-section can be described by the complex electric form factor $G_{E}$ and magnetic form factor $G_{M}$, which are functions of the four-momentum squared $q^{2}$ of the photon. The angular distribution of the produced baryon is characterized by two parameters: the angular parameter $\alpha_{0}$ and the phase difference between the EMFFs $\Delta\Phi$, where $\Delta\Phi$ is defined as $\text{arg}(G_{M})-\text{arg}(G_{E})$ [Chen:2019hqi]. The transverse polarization can be parameterized as
$\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}(\cos\theta_{0})=\frac{3}{2(3+\alpha_{0})}\sqrt{1-\alpha^{2}_{0}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi,$ (1)
where the factor $\frac{3}{2(3+\alpha_{0})}$ arises from normalization in spin density matrix of $\Lambda_{c}^{+}$. The $\alpha_{0}$ can be written as $\alpha_{0}=(|G_{E}|^{2}(1-v^{2})-|G_{M}|^{2})/(|G_{E}|^{2}(1-v^{2})+|G_{M}|^{2})$, where the $v$ is the velocity of the $\Lambda_{c}^{+}$ in center-of-mass (CM) system, normalized to the speed of light. A non-vanishing $\Delta\Phi$ gives rise to the transverse polarization effect of the baryon, which depends on the momentum direction of the baryon ($\theta_{0}$) in the rest frame of $e^{+}e^{-}$, as shown in Fig. 1.
Figure 1: Definition of the coordinate system and kinematic variables for the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$. The production angle $\theta_{0}$ is measured between the positron beam and the $\Lambda_{c}^{+}$ momentum in the center-of-mass frame, and the polarization $\mathcal{P}_{\Lambda_{c}^{+}}^{y}$ is oriented perpendicular to the production plane.
This polarization effect has been observed in the pair production of hyperons $\Lambda$ [BESIII:2018cnd, BESIII:2022qax, BESIII:2023euh], $\Sigma^{0,+}$ [BESIII:2023cvk, BESIII:2023ynq, BESIII:2024nif] and $\Xi^{0,-}$ [BESIII:2021ypr, BESIII:2022lsz, BESIII:2023drj] by the BESIII Collaboration, which has led to the highest precision measurements of the decay parameters [Lee:1957qs] of the hyperons. Similar to the pair production of hyperons, it is expected that the $\Lambda_{c}^{+}$ via $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ production can acquire polarization with respect to the normal of the production plane. However, this effect has never been established in charmed baryons.
Investigations of charmed baryon decay dynamics are essential for exploring the weak and strong interactions within the standard model (SM) of particle physics. The lightest charmed baryon, $\Lambda_{c}^{+}$, was discovered in 1979 [Abrams:1979iu]. It consists of one heavy charm quark and two light $u$ and $d$ quarks, being bound by the strong interaction. The charm quark mass lies in the intermediate region between perturbative and non-perturbative quantum chromodynamics (QCD), making the study of charmed baryons crucial for understanding non-perturbative effects in QCD.
In the weak decays of $\Lambda_{c}^{+}\to BP$, where $B$ and $P$ denote a spin-1/2 baryon and a pseudoscalar meson, respectively, the polarization of $\Lambda_{c}^{+}$, $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}$, can be transferred to the subsequent decays and manifested in the angular distributions of the daughter baryons, as shown in Fig. 2.
Figure 2: Generation and transfer of polarization in the weak decay $\Lambda_{c}^{+}\to BP$ [Wang:2024wrm]. The polarization of the daughter baryon $\bm{\mathcal{P}_{B}}$ depends on that of the parent $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}$, the decay angle $\theta$, and the weak decay parameters $\alpha$, $\beta$, and $\gamma$.
The polarization of the daughter baryon, denoted by $\bm{\mathcal{P}_{B}}$, is expressed as
$\small\bm{\mathcal{P}_{B}}=\frac{(\alpha+\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\cdot\bm{\hat{n}})\bm{\hat{n}}+\beta(\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\times\bm{\hat{n}})+\gamma\bm{\hat{n}}\times(\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\times\bm{\hat{n}})}{1+\alpha\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\cdot\bm{\hat{n}}}$, (2)
where $\alpha$, $\beta$, and $\gamma$ are the decay asymmetry parameters of $\Lambda_{c}^{+}\to BP$ and $\bm{\hat{n}}$ is the unit vector along the baryon flight direction in the rest frame of $\Lambda_{c}^{+}$. Previous study of the transverse polarization of the $\Lambda_{c}^{+}$ using data collected at a CM energy of $4600$ MeV at BESIII, only had a statistical significance of 2.1 standard deviations [BESIII:2019odb].
Various phenomenological models taking into account non-perturbative effects and non-factorizable contributions in QCD, have been developed to calculate the decay asymmetries of the charmed baryons. The development of these models can be roughly divided into three stages. The first stage was in the last century, after the CLEO Collaboration first measured $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ in 1990 [CLEO:1990unw]. Models such as covariant confined quark model (CCQM) [Korner:1992wi, Ivanov:1997ra], the pole model [Xu:1992vc, Cheng:1991sn, Cheng:1993gf, Zenczykowski:1993jm, Zenczykowski:1993hw], and current algebra (CA) [Datta:1995mn, Sharma:1998rd] were applied to calculate decay asymmetry parameters. The decay asymmetry parameters $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ calculated by different models varied significantly, particularly with regard to uncertainty on the sign. The beginning of the second stage is marked by the BESIII Collaboration’s measurement of the decay asymmetry in $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays together with the first determination of $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ in 2019 [BESIII:2019odb]. The central value of $\alpha_{pK^{0}_{S}}$ is positive with a large uncertainty, and most of the predictions were negative. Thus, the sign of $\alpha_{pK^{0}_{S}}$ remained a puzzle at that time. Another point is that the sign of $\alpha_{\Sigma^{0}\pi^{+}}$ has been confirmed. The Belle Collaboration results for $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ [Belle:2022uod, Belle:2022bsi] are in a good agreement with the flavor SU(3) symmetry. Subsequently, the BESIII Collaboration first reported the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay asymmetry parameters in 2024 [BESIII:2023wrw], which is a pure $W$-exchange process. The non-perturbative nature and the problem of non-factorization in this decay channel make calculations very difficult. A persistent tension exists between the branching fraction and the decay asymmetry parameters, as theoretical calculations and experimental measurements do not align perfectly. To resolve this tension, the study of the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay introduces a valuable perspective as the strong phase shift, which describes the relative phase difference induced by strong interactions in the decay, has a crucial role and may significantly influence the decay asymmetry parameters. Following this, some theoretical works [Geng:2023pkr, Zhong:2024zme, Zhong:2024qqs] have provided detailed discussions on this measurement and also introduced a strong phase shift, which significantly improves the description of the data. The most notable change due to the strong phase shift is seen for $\alpha_{pK^{0}_{S}}$, whose value becomes close to 0 [Zhong:2024qqs]. This modification leads to a substantial change in the related calculations, making the precise measurement of $\alpha_{pK^{0}_{S}}$ crucial. In 2024, the LHCb Collaboration measured $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Lambda\pi^{+}}$ with high accuracy using the $\Lambda_{b}^{0}\to\Lambda_{c}^{+}\pi^{-}$ decay, which provides a large longitudinal polarization to $\Lambda_{c}^{+}$ in the decay process  [LHCb:2024tnq]. These results agree with the latest theoretical calculations  [Cheng:2024lsn]. It is clear that the value of $\alpha_{pK^{0}_{S}}$ should be negative rather than close to zero. The BESIII Collaboration can also independently measure $\alpha_{pK^{0}_{S}}$ using a large sample to confirm the sign of $\alpha_{pK^{0}_{S}}$. For $\alpha_{\Lambda\pi^{+}}$, the Belle Collaboration reported $\alpha_{\Lambda\pi^{+}}$ [Belle:2022uod] with the same accuracy as the result from the LHCb Collaboration, but did not report $\beta_{\Lambda\pi^{+}}$ and $\gamma_{\Lambda\pi^{+}}$. All theoretical calculations and measurements about $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ are summarized in Table 1.
Table 1: Theoretical calculations and experimental measurements of the decay asymmetry parameters. The superscript $a$ means the model only considered SU(3) symmetry, and $b$ means considered the contribution of both SU(3) symmetry and SU(3) broken.
Calculations or experiments $\alpha_{pK^{0}_{S}}$ $\alpha_{\Lambda\pi^{+}}$ $\alpha_{\Sigma^{0}\pi^{+}}$ $\alpha_{\Sigma^{+}\pi^{0}}$
Körner(1992), CCQM [Korner:1992wi] 0.10-0.10 0.70-0.70 $0.70$ $0.71$
Xu(1992), Pole [Xu:1992vc] $0.51$ 0.67-0.67 $0.92$ $0.92$
Cheng, Tseng(1992), Pole [Cheng:1991sn] 0.49-0.49 0.96-0.96 $0.83$ $0.83$
Cheng, Tseng(1993), Pole [Cheng:1993gf] 0.49-0.49 0.95-0.95 $0.78$ $0.78$
Źencaykowski(1994), Pole [Zenczykowski:1993jm] 0.66-0.66 0.99-0.99 $0.39$ $0.39$
Źencaykowski(1994), Pole [Zenczykowski:1993hw] 0.90-0.90 0.86-0.86 0.76-0.76 0.76-0.76
Alakabha Datta(1995), CA [Datta:1995mn] 0.91-0.91 0.94-0.94 0.47-0.47 0.47-0.47
Ivanov(1998), CCQM [Ivanov:1997ra] 0.97-0.97 0.95-0.95 $0.43$ $0.43$
Sharma(1999), CA [Sharma:1998rd] 0.99-0.99 0.99-0.99 0.31-0.31 0.31-0.31
Geng(2019), SU(3) [Geng:2019xbo] $0.89^{+0.26}_{-0.11}$ $0.87\pm 0.10$ $0.35\pm 0.27$ $0.35\pm 0.27$
Zou(2020), CA [Zou:2019kzq] 0.75-0.75 0.93-0.93 0.76-0.76 0.76-0.76
Zhong (2022), SU(3)a [Zhong:2022exp] $0.51\pm 0.21$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Zhong (2022), SU(3)b [Zhong:2022exp] $0.29\pm 0.24$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Liu(2023), Pole [Liu:2023dvg] $0.81\pm 0.05$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Liu(2023), LP [Liu:2023dvg] $0.68\pm 0.01$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Geng(2023), SU(3) [Geng:2023pkr] $0.40\pm 0.49$ $0.75\pm 0.01$ $0.47\pm 0.02$ $0.47\pm 0.02$
Zhong(2024), TDA [Zhong:2024qqs] $0.01\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), IRA [Zhong:2024qqs] $0.03\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), TDA [Cheng:2024lsn] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
Zhong(2024), IRA [Cheng:2024lsn] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
CLEO(1990) [CLEO:1990unw] $1.0^{+0.4}_{-0.1}$
ARGUS(1992) [ARGUS:1991yzs] $0.96\pm 0.42$
CLEO(1995) [CLEO:1995qyd] $0.94^{+0.21+0.12}_{-0.06-0.06}$ $0.45\pm 0.31\pm 0.06$
FOCUS(2006) [FOCUS:2005vxq] $0.78\pm 0.16\pm 0.19$
BESIII(2019) [BESIII:2019odb] $0.18\pm 0.43\pm 0.14$ $0.80\pm 0.11\pm 0.02$ $0.73\pm 0.17\pm 0.07$ $0.57\pm 0.10\pm 0.07$
BELLE(2022) [Belle:2022uod, Belle:2022bsi] $0.755\pm 0.005\pm 0.003$ $0.463\pm 0.016\pm 0.008$ $0.48\pm 0.02\pm 0.02$
LHCb(2024) [LHCb:2024tnq] $0.754\pm 0.008\pm 0.006$ $0.785\pm 0.006\pm 0.003$
PDG Fit(2025) [ParticleDataGroup:2024cfk] $0.754\pm 0.010$ $0.768\pm 0.015$ $0.466\pm 0.018$ $0.484\pm 0.027$
The asymmetry in charge conjugation and parity symmetry (CPC\!P) lies at the heart of some of the most fundamental problems in physics. The presence of CPC\!P asymmetry has been established in meson decays [KTeV:1999kad, LHCb:2019hro, BaBar:2004gyj]. In 2019, the LHCb Collaboration observed a difference in time-integrated CPC\!P asymmetries between the decays $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ [LHCb:2019hro], providing the first observation of direct CPC\!P violation in the charm sector. This established that all three quark sectors, strange, bottom, and charm, exhibit CPC\!P-violating effects, albeit at very different magnitudes. Recently, the LHCb Collaboration reported the first observation of direct CPC\!P asymmetry in a baryonic decay [LHCb:2025ray], $\Lambda_{b}^{0}\to pK^{-}\pi^{+}\pi^{-}$, marking a significant milestone in the study of CPC\!P violation. Meanwhile, ongoing efforts are exploring CPC\!P asymmetry in other baryonic processes [LHCb:2016yco, LHCb:2024yzj]. Despite these discoveries, the origin of CPC\!P asymmetry and a precise determination of its magnitude in specific processes remain open questions. In the SM, CPC\!P violation arises from the interplay between strong and weak phases: weak phases change sign under a CPC\!P transformation, while strong phases remain invariant. A measurement of the CPC\!P asymmetry thus enables extraction of these two phases, and if one is known, the other can be determined. The strong phase shift is an important input for theoretical calculations, helping to refine models of hadronic dynamics and improve our understanding of the underlying strong interaction. On the other hand, the weak phase provides a direct test of SM predictions, offering valuable insights into potential CPC\!P asymmetry and verifying the consistency of weak interactions in charm decays. In Cabibbo-favored (CF) decays of the $\Lambda_{c}^{+}$, there is no CPC\!P asymmetry in SM, as the weak phase difference is expected to be zero. Although observing a non-zero weak phase shift in these processes is currently unrealistic based on the current data sample, a well-measured strong phase shift would still enhance the potential observability of a weak phase shift.
In this paper, we observe for the first time the transverse polarization of the $\Lambda_{c}^{+}$ baryon in $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ and obtain the phase difference ($\Delta\Phi$) between the EMFFs at thirteen CM energies ranging from $4600$ to $4951$ MeV, with an overall significance exceeding 10.0 standard deviations. The decay asymmetry parameters ($\alpha_{BP}$ and $\Delta_{BP}$, which then leads to $\beta_{BP}$ and $\gamma_{BP}$) and phase shifts between the $S$- and $P$-wave ($\delta_{p}-\delta_{s}$) are measured for four CF channels: $\Lambda_{c}^{+}\to pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$. The decay of $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, notable for its clean structure and large branching fraction, is included to improve the transverse polarization measurement. Unless otherwise specified, the charge-conjugate channels are included throughout this article. By measuring the processes in samples separated by baryon charge, the CPC\!P parameters, strong phase shifts and weak phase shifts between different partial waves are also extracted in the CF processes. This allows precise testing of the SM and provides an essential input for the theoretical calculations. The phase differences $\delta_{p}-\delta_{s}$ and $\Delta\Phi$ encode the characteristics of the $\Lambda_{c}^{+}$ decay and production processes, respectively. Besides the involved strong interaction, the former is associated with the weak interaction, while the latter connects the electromagnetic one. CPC\!P violation can arise only from the weak interaction accompanying with the strong interaction. The analysis is performed based on a multidimensional angular analysis of the cascade decays of $\Lambda_{c}^{+}$, as described in Ref. [BESIII:2019odb].
II Experimental apparatus and data sets
The Beijing Spectrometer III (BESIII) detector [BESIII:2009fln] records symmetric $e^{+}e^{-}$ collisions provided by the Beijing Electron Positron Collider (BEPCII) storage ring [Yu:2016cof] in the CM energy range from 1.84 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.773\;\text{GeV}$. BESIII has collected large data samples in this energy region [BESIII:2020nme, Lu2020-hf, Zhang2022-id]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps [Guo2017-of, Li2017-zt, Cao2020-vl]. Data analyzed in this work were collected at thirteen energy points above the $\Lambda^{+}_{c}\bar{\Lambda}^{-}_{c}$ threshold. The integrated luminosity of the datasets analyzed in this work is estimated to be $6.4~\mathrm{fb}^{-1}$ [BESIII:2022dxl, BESIII:2022ulv].
Simulated data samples produced with a geant4-based [GEANT4:2002zbu] Monte Carlo package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The simulation models the beam energy spread and initial state radiation (ISR) in the $e^{+}e^{-}$ annihilations with the generator kkmc [Jadach:2000ir, Jadach:1999vf]. The inclusive simulation includes the production of open charm processes, the ISR production of vector charmonium(-like) states, and the continuum processes incorporated in kkmc [Jadach:2000ir, Jadach:1999vf]. All particle decays are modelled with evtgen [Lange:2001uf, Ping:2008zz] using branching fractions either taken from the Particle Data Group (PDG) [ParticleDataGroup:2024cfk], when available, or otherwise estimated with lundcharm [Chen:2000tv, Yang:2014vra]. Final state radiation (FSR) from charged final state particles is incorporated using the photos package [Barberio:1990ms]. The phase space (PHSP) signal simulation is generated according to a pure phase-space model, assuming a uniform distribution of the available kinematic phase space without any dynamical amplitudes or polarization effects for the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, followed by the decay of $\Lambda_{c}^{+}$ into the signal final states and the inclusive decay of $\bar{\Lambda}_{c}^{-}$, which accounts for all possible decay modes. The signal simulation is based on the process model described in the helicity formalism with the decay asymmetry parameters measured in this work or cited from the other measurements [BESIII:2022qax, ParticleDataGroup:2024cfk].
III Measurement Method
III.1 Event selection and $\Lambda_{c}^{+}$ signal
in the beam-constrained mass spectrum
A single tag approach [BESIII:2017kqg] is employed, in which an event is accepted as an exclusive $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ final state when either a $\Lambda_{c}^{+}$ or its antiparticle is reconstructed in one of the decay channels studied, using all particles in the same side as the reconstructed baryon. Four two-body hadronic decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, $\Sigma^{+}\pi^{0}$, and one three-body decay, $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, are studied in this research following Refs. [BESIII:2017kqg, BESIII:2022xne]. The intermediate states, $K_{S}^{0}$, $\Lambda$, $\Sigma^{0}$, and $\Sigma^{+}$ are reconstructed by combining $\pi^{+}\pi^{-}$, $p\pi^{-}$, $\gamma\Lambda$, and $p\pi^{0}$ candidates, respectively.
Charged tracks detected in the MDC are required to satisfy $|\cos\theta|<0.93$, where $\theta$ is the polar angle defined relative to the symmetry axis of the MDC (the $z$-axis). Additionally, the tracks that do not come from a $K^{0}_{S}$ meson or $\Lambda$ baryon must have a distance of closest approach to the interaction point (IP) satisfying $|r_{z}|<10~\text{cm},\quad|r_{xy}|<1~\text{cm}$, where $r_{z}$ and $r_{xy}$ are the distances along the beam axis and in the transverse plane, respectively.
Particle identification (PID) for charged tracks combines measurements of the energy deposited in the MDC (dEE/dxx) and the flight time in the TOF to form likelihoods $\mathcal{L}(h)~(h=p,K,\pi)$ for each hadron $h$ hypothesis. Tracks are identified as protons if their likelihoods satisfy $\mathcal{L}(p)>\mathcal{L}(K)$ and $\mathcal{L}(p)>\mathcal{L}(\pi)$. Charged kaons and pions are identified by comparing the likelihoods of the respective hypotheses: $\mathcal{L}(K)>\mathcal{L}(\pi)$ and $\mathcal{L}(\pi)>\mathcal{L}(K)$, respectively.
Neutral showers are reconstructed in the EMC. Showers not associated with any charged track are identified as photon candidates. The deposited energy of each shower in the EMC must satisfy the following conditions: $E_{\text{barrel}}>25$ MeV for $|\cos\theta|<0.80$, $E_{\text{end-cap}}>50$ MeV for $0.86<|\cos\theta|<0.92$. The EMC time difference from the event start time is required to be less than 700 ns to suppress electronic noise and showers unrelated to the event, where the event start time corresponds to the time at which the $e^{+}e^{-}$ collision occurs. The $\pi^{0}$ candidates are reconstructed from photon pairs with invariant masses satisfying: $115~\text{MeV}/c^{2}<M(\gamma\gamma)<150~\text{MeV}/c^{2}$.
Candidates for $K^{0}_{S}$ and $\Lambda$ decays are reconstructed from $\pi^{+}\pi^{-}$ and $p\pi^{-}$ combinations, respectively. For these tracks, the distances of closest approach to the interaction point (IP) must satisfy: $|r_{z}|<20~\text{cm}$. There is no distance constraint required in the transverse plane. The charged pions in the $K^{0}_{S}$ reconstruction are not subject to the PID requirements described earlier, while the proton in the $\Lambda$ reconstruction is required to satisfy the PID criteria. The two final-state tracks are constrained to originate from a common decay vertex by requiring the vertex fit to satisfy: $\chi^{2}<100$. Furthermore, the decay vertex is required to be separated from the IP by a distance of at least twice the fitted vertex resolution. To select $K^{0}_{S}$ and $\Lambda$ candidates, the invariant masses of the combinations must satisfy: $487~\text{MeV}/c^{2}<M(\pi^{+}\pi^{-})<511~\text{MeV}/c^{2}$, $1111~\text{MeV}/c^{2}<M(p\pi^{-})<1121~\text{MeV}/c^{2}$. These mass windows correspond to approximately three times the standard deviation on either side of the known masses.
To suppress the peaking backgrounds for some decay channels, like $\Lambda_{c}^{+}\to\Lambda l^{+}\nu_{l}(l=e/\mu)$ background in the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, and $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, additional requirements are applied. Positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. For the $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ channel, it is required that $\mathcal{L}^{\prime}(e)<0.05$ and $\mathcal{L}(\pi)>0.001$ are satisfied to suppress the $\Lambda_{c}^{+}\to\Lambda e^{+}\nu_{e}$ background, which rejects roughly 95% of these backgrounds with only a negligible signal loss. For the $\Lambda_{c}^{+}\to\Lambda\mu^{+}\nu_{\mu}$ background in the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, the information from the muon counter detector is utilized, and the penetrated depth of the $\pi^{+}$ candidates is required to be less than $5~\text{cm}$. These selections effectively reduce peaking backgrounds $\Lambda_{c}^{+}\to\Lambda\mu^{+}\nu_{\mu}$ by 80% while keeping over 80% of the signal events. A kinematic variable, the energy difference $\Delta E\equiv E_{\Lambda_{c}^{+}}-E_{\text{beam}}$, is used to suppress background, where $E_{\Lambda_{c}^{+}}$ is the total measured energy of the $\Lambda_{c}^{+}$ candidate, and $E_{\text{beam}}$ is the beam energy. For the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$, $pK^{0}_{S}$, $\Lambda\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays the requirements are $\Delta E\in(-29,26)~\text{MeV}$, $\Delta E\in(-21,18)~\text{MeV}$, $\Delta E\in(-23,21)~\text{MeV}$, and $\Delta E\in(-67,32)~\text{MeV}$, respectively. For the $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ channel, an additional photon can be selected from the process $\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and its $\Delta E$ is expected to be larger than zero. The requirement $\Delta E\in(-33,20)~\text{MeV}$ removes about 54% of this type of backgrounds while retaining more than 98% of signal events.
A large background for the channel $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$ is from $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ combined with a random photon. This background, which may pose hidden dangers to the angular analysis, necessitates the application of a four-constraint kinematic fit to the $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$, where $\Sigma^{+}$ decays to proton and $\pi^{0}$, $\pi^{0}\to\gamma\gamma$. The quality of the fit, $\chi^{2}$, is calculated by constraining two $\pi^{0}$s, $\Sigma^{+}$, and $\Lambda_{c}^{+}$ to their world average masses quoted from PDG [ParticleDataGroup:2024cfk]. After optimization by $N^{2}_{S}/(N_{S}+N_{B})^{3/2}$ [LHCb:2021chn, LHCb:2022ogu], the requirement is set that the $\chi^{2}$ must be less than 12, where $N_{S}$ and $N_{B}$ represent the signal and background yields, respectively.
After applying the above selection criteria, the beam-constrained mass $M_{\text{BC}}\equiv\sqrt{E_{\text{beam}}^{2}/c^{4}-|\vec{p}_{\Lambda_{c}^{+}}|^{2}/c^{2}}$, is defined to identify signal candidates, and $\vec{p}_{\Lambda_{c}^{+}}$ is the momentum of the $\Lambda_{c}^{+}$ candidates calculated in the $e^{+}e^{-}$ rest frame. An unbinned maximum-likelihood fit is performed on the distribution of $M_{\text{BC}}$. The shapes of both correctly and mis-reconstructed signal events are modeled with the simulation, convolved with a Gaussian function to account for the resolution difference between the data and the simulation. The truth matching technique classifies events as correctly reconstructed if the angle between the reconstructed and true photon momentum directions from the simulation is below 10 degrees; otherwise, they are mis-reconstructed. The background shapes are described by an ARGUS function [ARGUS:1990hfq], except in the case of the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channel at CM energies larger than $4740$ MeV, where a second-order Chebyshev polynomial function is employed to model the background. The fit results are shown in Fig. 3 for all channels at $4600$ MeV. The obtained signal yields, combining all energy points, are $50083\pm 233$, $9619\pm 96$, $5742\pm 74$, $2487\pm 48$ and $1268\pm 37$ in the $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$$pK^{0}_{S},~\Lambda\pi^{+},~\Sigma^{0}\pi^{+}\text{,~and}~\Sigma^{+}\pi^{0}$ channels, respectively, within the signal region of $M_{\text{BC}}\in[2.282,2.291]~\text{GeV}/c^{2}$.
Figure 3: Beam-constrained mass spectrum for $\Lambda_{c}^{+}$ signal at $4600$ MeV. Each panel shows a different signal channel as labeled. The Black points with error bars are data. The black shaded region indicates the mis-reconstructed signal events. The green region represents the fitted combinatorial background.
III.2 Helicity system and angular distribution definitions
A helicity system and helicity angles are employed in constructing the amplitude for $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, $\Lambda_{c}^{+}\to BP$. Taking the channel $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ as an example, a kinematic diagram of this decay is shown in Fig. 4. For the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, the momenta of the $\Lambda_{c}^{+}$ and $\bar{\Lambda}_{c}^{-}$ baryons are defined in the CM system of the $e^{+}e^{-}$ pair, and $\theta_{0}$ is the polar angle of the $\Lambda_{c}^{+}$ in CM system. For the helicity system describing $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ decay, the momenta of the $\Lambda$ and the $\pi^{+}$ are defined in the $\Lambda_{c}^{+}$ frame. The angle $\phi_{1}$ is that between the $e^{+}\Lambda_{c}^{+}$ and $\Lambda\pi^{+}$ planes, and $\theta_{1}$ is the polar angle of the $\Lambda$ momentum defined in the rest frame of the $\Lambda_{c}^{+}$ baryon. For the helicity system describing the $\Lambda\to p\pi^{-}$, $\phi_{2}$ and $\theta_{2}$ adopt similar definitions. Then the formula of angular distributions related to this helicity frame based on the language of helicity amplitudes and Wigner-D function can be derived [BESIII:2019odb].
Figure 4: Definitions of the helicity frames and related angles for $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-},\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and $\Lambda\to p\pi^{-}$.
The joint angular distribution of the single-side decay chains can be derived from polarization transmission [Chen:2019hqi] or equivalently the helicity density matrix [Perotti:2018wxm]. The angular distribution of the decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to\Lambda(\to p\pi^{-})\pi^{+})$ is shown below:
$\small\begin{split}&\frac{\text{d}\varGamma}{\text{d}\cos\theta_{0}~\text{d}\cos\theta_{1}~\text{d}\cos\theta_{2}~\text{d}\phi_{1}~\text{d}\phi_{2}}\\ &\propto 1+\alpha_{0}\cos^{2}\theta_{0}\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{p\pi^{-}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\sin\theta_{1}\sin\phi_{1}\cos\theta_{2}\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{p\pi^{-}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\cos\theta_{1}\sin\phi_{1}\sin\theta_{2}\\ &\times\sqrt{1-\alpha^{2}_{\Lambda\pi^{+}}}\cos(\Delta_{\Lambda\pi^{+}}+\phi_{2})\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{p\pi^{-}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\cos\phi_{1}\sin\theta_{2}\\ &\times\sqrt{1-\alpha^{2}_{\Lambda\pi^{+}}}\sin(\Delta_{\Lambda\pi^{+}}+\phi_{2})\\ &+\sqrt{1-\alpha^{2}_{0}}\alpha_{\Lambda\pi^{+}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\sin\theta_{1}\sin\phi_{1}\\ &+\alpha_{0}\alpha_{p\pi^{-}}\alpha_{\Lambda\pi^{+}}\cos^{2}\theta_{0}\cos\theta_{2}\\ &+\alpha_{p\pi^{-}}\alpha_{\Lambda\pi^{+}}\cos\theta_{2},\end{split}$ (3)
where the $\alpha_{p\pi^{-}}$ denotes the decay asymmetry parameter in the weak hadronic decay $\Lambda\to p\pi^{-}$. The parameter $\Delta_{\Lambda\pi^{+}}$ is the phase shift between the helicity amplitudes of $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ satisfying $\beta_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\sin\Delta_{\Lambda\pi^{+}}$ and $\gamma_{\Lambda\pi^{+}}=\sqrt{1-(\alpha_{\Lambda\pi^{+}})^{2}}\cos\Delta_{\Lambda\pi^{+}}$. Formulas of other two-body decays can be found in Ref. [BESIII:2019odb].
Although the analysis of the angular distribution of these four two-body CF channels provides effective constraints, the three-body channel $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, with a low background and large production rate, brings beneficial additional sensitivity. Benefiting from the transition amplitude analysis reported by the LHCb Collaboration [LHCb:2023crj], the angular distribution of decay chain $e^{+}e^{-}\to\gamma^{*}\to\bar{\Lambda}_{c}^{-}\Lambda_{c}^{+}(\to pK^{-}\pi^{+})$ can be expressed explicitly as:
$\small\begin{split}&\frac{\text{d}\mathcal{\varGamma}}{\text{d}\cos\theta_{0}~\text{d}\cos{\theta_{1}}~\text{d}\phi_{1}~\text{d}\phi_{2}}\\ &\propto I_{0}(\kappa)\\ &+I_{0}(\kappa)\cos^{2}\theta_{0}\\ &+I_{0}(\kappa)\sqrt{1-\alpha^{2}_{0}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi\sum_{j}R_{2j}(\phi_{1},\theta_{1},\phi_{2})\alpha_{j}(\kappa),\end{split}$ (4)
where $\phi_{1},\theta_{1},\phi_{2}$ are the helicity angles for the intermediate state ($K^{-}\pi^{+}$), $R_{ij}(\phi_{1},\theta_{1},\phi_{2})$ represents the 3-dimension $Z$-$Y$-$Z$ rotation matrix and $\kappa$ denotes a set of Dalitz-plot variables $(M(pK^{-})^{2},M(K^{-}\pi^{+})^{2})$. The $\vec{\alpha}(\kappa)$ and $I_{0}(\kappa)$ are model-agnostic representations for polarization dependence of the decay rate and the total differential decay rate over $\kappa$, respectively, which are taken from the $\Lambda_{c}^{+}$ polarimetry with the amplitude model both measured by the LHCb Collaboration [LHCb:2023crj, LHCb:2022sck].
III.3 Fit of the joint angular distributions
In this analysis, the free parameters of $\alpha_{0}$, $\Delta\Phi$, $\alpha_{BP}$, and $\Delta_{BP}$, which describe the angular distributions for each CM energy are determined through a simultaneous unbinned maximum log-likelihood fit. The $\alpha_{0}$ and $\Delta\Phi$ are specific to the invididual energy point. The likelihood function is constructed from the joint probability density function (PDF) by
$\mathcal{L}_{\text{total}}=\sum^{\text{energy}}\mathcal{L}_{\text{data}}=\prod_{i=1}^{N_{\text{data}}}f_{s}(\vec{\xi}_{i}).$ (5)
Here, $f_{s}(\vec{\xi}_{i})$ is the PDF of the signal process, $N_{\text{data}}$ is the number of data events and $i$ is the event index. The signal PDF $f_{s}(\vec{\xi}_{i})$ is formulated as
$f_{s}(\vec{\xi}_{i})=\frac{\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}}{\int\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}\text{d}\vec{\xi}_{i}}~,$ (6)
where $\vec{\xi}_{i}$ denotes the kinematic observables, including the helicity angles and Dalitz-plot variables, while $\vec{\eta}$ denotes the free parameters to be determined. The $M(\vec{\xi}_{i};\vec{\eta})$ is the total decay amplitude, and $\epsilon(\vec{\xi}_{i})$ is the detection efficiency, parameterized in terms of the kinematic variables $\vec{\xi}_{i}$. It is important to note that all $\vec{\xi}_{i}$ are calculated using four-momenta updated by a kinematic fit, which constrains the invariant mass of all resonances in the signal decay to their world average masses [ParticleDataGroup:2024cfk]. The integration of the normalization factor is calculated with a large PHSP signal simulation samples as $\int\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}\text{d}\vec{\xi}_{i}=\frac{1}{N_{\text{gen}}}\sum_{k_{\text{surv}}}^{N_{\text{surv}}}|M(\vec{\xi}_{k_{\text{surv}}};\vec{\eta})|^{2}$, where $N_{\text{gen}}$ is the total number of the generated events, $N_{\text{surv}}$ is the number of the events surviving all selection criteria and $k_{\text{surv}}$ is the event index. The background contribution to the joint likelihood is subtracted based on the calculated likelihoods using data in $M_{\text{BC}}\in[2.250,2.270]~\text{GeV}/c^{2}$. For the mis-reconstructed signal events, the contribution to the likelihood is based on the signal MC simulation. Both components are scaled to match the corresponding yields in data, and the scale factors can be obtained from the $M_{\text{BC}}$ fitting. The likelihood function with background subtraction can be expressed as
$\mathcal{L}_{\text{sig}}=K\times(\mathcal{L}_{\text{data}}-\mathcal{L}_{\text{bkg}}-\mathcal{L}_{\text{mis}}),$ (7)
where the factor $K$ is used to correct uncertainties in unbinned maximum likelihood fit as suggested in Refs. [BESIII:2022udq, Langenbruch:2019nwe]. The subscripts “bkg” and “mis” correspond to the background and mis-reconstructed components, respectively.
The negative log-likelihood with background subtraction over the five signal channels and data samples collected from thirteen energy points is minimized using the Minuit [James:1975dr] algorithms. The decay asymmetry parameters of the weak decays $\Lambda\to p\pi^{-}$ and $\Sigma^{+}\to p\pi^{0}$ and their charge conjugated processes $\bar{\Lambda}\to\bar{p}\pi^{+}$ and $\bar{\Sigma}^{-}\to\bar{p}\pi^{0}$ are fixed according to the experimentally measured decay asymmetry parameters [BESIII:2022qax, ParticleDataGroup:2024cfk]. In the nominal fit, the parameters $\alpha_{BP}$ and $\Delta_{BP}$ are set to $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$, respectively, in accordance with the assumption of CPC\!P conservation. Here, $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ represent the parameters associated with the antiparticle $\bar{\Lambda}_{c}^{-}$ decays. The results of the fit are presented in Fig. 5, and the full correlation matrix is provided in Table A1 of Appendix A. An alternative test is also performed where the assumption of CPC\!P conservation are relaxed and $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ are treated as free parameters in the fit.
Figure 5: Fit results of helicity angles combining all energy points. They are angular distribution about $pK_{S}^{0}$, $pK^{-}\pi^{+}$, $\Lambda\pi^{+}$, $\Sigma^{+}\pi^{0}$, and $\Sigma^{0}\pi^{+}$, respectively. Black points with error bars are data; the green shaded region indicates combinatorial background events and black shaded region is the mis-reconstructed signal events.
III.4 Systematic uncertainties
The systematic uncertainties originate from the differences between simulations and real physical processes leading to deviations in detection efficiencies. In this part, the absolute systematic uncertainties are estimated. The numerical values are summarized in Table 2, and the individual sources are discussed in the rest of this section.
[…]
  • Performance of truth matching. In the nominal measurement, the angle between the reconstructed and true photon momentum directions from simulations are required to be below 10 degrees. To evaluate the effect due to this truth matching requirement, two different matching conditions of 9 and 11 degrees are tested and the resultant largest difference from the nominal result is considered as systematic uncertainty.
  • Signal components. The amplitude-model related uncertainties are computed using the alternative models in Ref. [LHCb:2022sck]. With alternative $\Lambda^{+}_{c}$ polarimetry, the angular fit is repeated and the standard deviations and difference of mean value are assigned to the model uncertainties. In Ref. [LHCb:2023crj], the statistical uncertainties are estimated by varying the parameter values in the default model. These parameters are sampled from a Gaussian function, whose $\mu$ and $\sigma$ are given by the central and error values from the default model fit results in Ref. [LHCb:2022sck]. The polarimeter vector fields are reproduced for each varied parameter set. The angular fit procedure is repeated using the corresponding inputs, and the statistical component is given by taking the standard deviations and the difference of the mean value.
  • Background components. Background fractions are fixed according to the fit results from the nominal fit. To estimate systematic uncertainties, the background fraction for each energy point is varied by $\pm 1\sigma$ and the fit procedure is repeated. The quadratic sum of the largest variation from each fraction are assigned as systematic uncertainties from the background fraction. The mis-reconstructed model is examined using an alternative signal simulation produced with new input parameters $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ which are changed within $\pm 1\sigma$ of the results presented here. The background shape is extracted from the data events in the $M_{\text{BC}}$ sideband regions. For $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channel, to estimate the possible difference between the $M_{\text{BC}}$ sideband and signal regions, a weight factor is derived by comparing the $M(K^{-}\pi^{+})^{2}$ distributions in the $M_{\text{BC}}$ sideband and signal regions of inclusive simulation, since the most significant discrepancy appears in the $M(K^{-}\pi^{+})^{2}$ distribution, while the others show consistent behavior. Data events in the $M_{\text{BC}}$ sideband region are corrected by the weight factor and the fit procedure is repeated. The difference is then taken as the uncertainty from background model. For two-body decays, the uncertainty due to the combinational background model is estimated by varying the relative weights between $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs and other hadronic events based on the uncertainties of their cross-section ratio. The relevant systematic uncertainties from background size are examined by repeating the fits with an alternative background size obtained from the Gaussian sampling of the fitted parameters. The series of fitted parameters, assumed to follow a Gaussian distribution, are used to estimate the systematic uncertainty for the background size. The fitted Gaussian means are found to be consistent to the nominal result, typically with central value deviating at the level of $10^{-3}$. To be conservative, the corresponding systematic uncertainty is determined by summing the fitted Gaussian width and the deviation of the fitted Gaussian mean from the nominal value.
  • Fixed parameters. The systematic uncertainty due to the input parameters is evaluated by varying these parameters using a Gaussian sampling method. For each parameter, the obtained results are expected to follow Gaussian distributions. The sum of the fitted Gaussian resolution and the difference between the fitted Gaussian mean and the nominal result is taken as the systematic uncertainty. The fitted means are close to zero [Zhong:2024qqs], but they are included with the width as a conservative estimation.
Table 2: Absolute systematic uncertainties in transverse polarization and decay asymmetry parameters (in percent), where SEL, TMP, SIG, BKG, and PAR in the table represents the sources from Selection criteria, Performance of truth matching, Signal components, Background components, and Fixed parameters, respectively. Some negligible terms are denoted by “”. The $\alpha_{0}$ and $\Delta\Phi$ are given as ranges because they vary for different energy points. The total uncertainty is calculated as the sum in quadrature of all components.
Parameters SEL TMP SIG BKG PAR Total
$\alpha_{0}$ $\leq 0.2$ $\leq 0.1$ $\leq 0.1$ 0.11.90.1-1.9 $\leq 0.1$ 0.11.90.1-1.9
$\Delta\Phi$ $\leq 0.3$ $\leq 0.1$ 0.45.80.4-5.8 0.62.50.6-2.5 $\leq 0.6$ 0.95.80.9-5.8
$\alpha_{pK^{0}_{S}}$ $0.1$ $1.1$ $2.9$ $0.1$ $3.1$
$\alpha_{\Lambda\pi^{+}}$ $0.8$ $0.3$ $0.3$ $0.9$
$\Delta_{\Lambda\pi^{+}}$ $0.9$ $0.8$ $0.8$ $0.1$ $1.4$
$\alpha_{\Sigma^{0}\pi^{+}}$ $0.3$ $0.1$ $0.8$ $0.2$ $0.9$
$\Delta_{\Sigma^{0}\pi^{+}}$ $1.5$ $1.0$ $2.2$ $0.3$ $2.9$
$\alpha_{\Sigma^{+}\pi^{0}}$ $1.1$ $1.2$ $1.4$ $2.2$
$\Delta_{\Sigma^{+}\pi^{0}}$ $1.2$ $1.2$ $3.7$ $0.3$ $4.0$
IV Results
IV.1 Transverse polarization
The results of the angular distribution fit, the $\alpha_{0}$, the phase difference $\Delta\Phi$, its significance, as well as the largest transverse polarization $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y,max}}$, are summarized in Table 3. The EMFF-related $\alpha_{0}$ is consistent with previous results but more precise. The ratios $|G_{E}/G_{M}|$ are compared with the previous BESIII measurements [Chen:2023oqs] and theoretical calculations [Wan:2021ncg, Chen:2023oqs], as shown in Fig. 6. The measurements of $\sin\Delta\Phi$ are also presented with the theoretical calculations [Wan:2021ncg, Chen:2023oqs], which significantly differs from the theoretical predictions, providing information beyond current theoretical expectations.
Table 3: Results for $\alpha_{0}$, $\Delta\Phi$, and transverse polarization of $\Lambda_{c}^{+}$ at 13 energy points, where the first uncertainties are statistical and the second are systematic. In the table, SLSL means the significance level of non-zero $\Delta\Phi$, $\sigma$ represents the standard deviation, and $\bm{{\mathcal{P}_{\Lambda_{c}^{+}}^{y,max}}}$ is the largest transverse polarization.
$\sqrt{s}~[\text{MeV}]$ $\alpha_{0}$~in previous work [BESIII:2019odb, BESIII:2023rwv] $\alpha_{0}$~in this work $\Delta\Phi~[\text{rad}]$ SLSL $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y,max}}$
$4600$ $0.20\pm 0.04\pm 0.02$ $0.226\pm 0.030\pm 0.004$ $0.100\pm 0.069\pm 0.009$ $2.2~\sigma$ $0.026\pm 0.018\pm 0.002$
$4612$ $0.26\pm 0.09\pm 0.01$ $0.160\pm 0.083\pm 0.004$ $0.146\pm 0.162\pm 0.030$ $1.1~\sigma$ $0.038\pm 0.042\pm 0.008$
$4628$ $0.21\pm 0.04\pm 0.01$ $0.181\pm 0.038\pm 0.001$ $0.371\pm 0.082\pm 0.012$ $6.8~\sigma$ $0.095\pm 0.020\pm 0.003$
$4641$ $09\pm 0.05\pm 0.01 0.0$$060\pm 0.039\pm 0.003$ $0.398\pm 0.073\pm 0.015$ $7.6~\sigma$ $0.099\pm 0.017\pm 0.004$
$4661$ $02\pm 0.05\pm 0.01 0.0$$008\pm 0.044\pm 0.003$ $0.496\pm 0.088\pm 0.021$ $8.5~\sigma$ $0.119\pm 0.019\pm 0.005$
$4682$ $0.15\pm 0.03\pm 0.01$ $0.102\pm 0.029\pm 0.003$ $0.502\pm 0.054\pm 0.021$ $14.1~\sigma$ $0.116\pm 0.011\pm 0.005$
$4699$ $0.34\pm 0.07\pm 0.01$ $0.305\pm 0.055\pm 0.010$ $0.545\pm 0.114\pm 0.028$ $7.1~\sigma$ $0.112\pm 0.021\pm 0.007$
$4740$ $0.49\pm 0.16\pm 0.03$ $358\pm 0.126\pm 0.008 0.0$$097\pm 0.190\pm 0.016$ $4~\sigma 0.0$$020\pm 0.039\pm 0.004$
$4750$ $0.42\pm 0.10\pm 0.01$ $0.347\pm 0.079\pm 0.004$ $0.316\pm 0.142\pm 0.019$ $3.1~\sigma$ $0.065\pm 0.029\pm 0.005$
$4781$ $0.17\pm 0.07\pm 0.01$ $0.157\pm 0.062\pm 0.007$ $0.395\pm 0.126\pm 0.028$ $5.1~\sigma$ $0.090\pm 0.027\pm 0.007$
$4843$ $0.38\pm 0.10\pm 0.01$ $0.282\pm 0.089\pm 0.019$ $0.385\pm 0.153\pm 0.034$ $3.6~\sigma$ $0.082\pm 0.031\pm 0.008$
$4918$ $0.62\pm 0.17\pm 0.01$ $0.612\pm 0.150\pm 0.019$ $0.423\pm 0.272\pm 0.024$ $1.9~\sigma$ $0.067\pm 0.043\pm 0.013$
$4951$ $0.63\pm 0.21\pm 0.01$ $0.744\pm 0.179\pm 0.007$ $0.700\pm 0.392\pm 0.058$ $1.8~\sigma$ $0.086\pm 0.050\pm 0.030$
Figure 6: Comparison of $|G_{E}/G_{M}|$ and $\sin\Delta\Phi$ between theoretical calculation and this work. Red points with error bars correspond to this work, and the gray ones are previous measurement results from BESIII [BESIII:2019odb, BESIII:2023rwv]. The purple dashed line with sky blue shadow is calculated from Ref. [Wan:2021ncg], and the blue solid line represents the calculated results with Ref. [Chen:2023oqs].
Of particular interest is the fact that the $\Lambda_{c}^{+}$ baryon is significantly polarized with respect to the normal of the production plane. The size of polarization is strongly dependent on $\Lambda_{c}^{+}$ momentum direction $\cos\theta_{0}$ with a maximum value of approximately 12%. To illustrate the polarization effect, the moment $\textless\sin\theta_{1}\sin\phi_{1}\textgreater$ is calculated as the average of $\sin\theta_{1}\sin\phi_{1}$ within a certain range of $\cos\theta_{0}$. It is expected to be proportional to $\alpha_{BP}P_{y}(\cos\theta_{0})/(1+\alpha_{0}\cos^{2}\theta_{0})$. We divide the range of $\cos\theta_{0}$, (1,1)(-1,1), into eight bins, and calculate the moment for experimental data and provide the best fit curve at all energy points, as shown in Fig. 7. To compute the significance, we perform a floating fit for $\Delta\Phi$ and a fixed fit for $\Delta\Phi=0$ to obtain the likelihood values and the degrees of freedom, from which we derive the significance of non-zero $\Delta\Phi$. The sign($\alpha_{BP}$) is used to avoid cancellation between charge-conjugated processes. The significance is reported in Table 3.
Figure 7: The moment $\textless\text{sign}(\alpha_{BP})\sin\theta_{1}\sin\phi_{1}\textgreater$ distribution to show the effect of transverse polarization. Points with error bars correspond to data with background subtracted and the red solid line denote the global expected dependence on $\cos\theta_{0}$. The blue dashed line represents the distribution without polarization uniformly distributed in phase space.
IV.2 Decay asymmetry and SM test
As the SM predicts no CPC\!P asymmetry in these decay channels, we assume CPC\!P symmetry, meaning the decay asymmetry parameters $\alpha_{BP}=-\bar{\alpha}_{BP}$ and $\Delta_{BP}=-\bar{\Delta}_{BP}$ in the angular distribution analysis. In addition to the transverse polarization, the decay asymmetry parameters are obtained and shown in Fig. 8. In the SM, the amplitude for a spin-$1/2$ baryon decaying into a spin-$1/2$ baryon and a spin-0 meson can be written as $\mathcal{M}=i\bar{u}_{f}(A-B\gamma_{5})u_{i}$, where $A$ and $B$ are constants, that reflect the $S$ and $P$-wave amplitudes with different orbital angular momentum in weak decays, arising from the parity-violating and parity-conserving components of the weak interaction, respectively. The $u_{i}$ and $\bar{u}_{f}$ are spinors describing the initial and final baryons [Xu:1992vc]. By combining the measured parameter $\alpha_{BP}$, $\Delta_{BP}$, and the branching fraction $\mathcal{B}$, a set of equations can be established to relate the observables to the decay amplitudes. These equations allow the extraction of the magnitudes of $|A|$ and $|B|$ as well as the relative phase shift $\delta_{p}-\delta_{s}$:
[…]
where $\vec{p}_{c}$ is the momentum of $B$ in the rest frame of $\Lambda_{c}^{+}$ and the auxiliary parameter $\kappa$ is defined as $\kappa=|\vec{p}_{c}|/(E_{B}+m_{B})=\sqrt{(E_{B}-m_{B})/(E_{B}+m_{B})}$. The symbol $\tau_{\Lambda_{c}^{+}}$ denotes the mean lifetime of the baryon, and $E$, $m$ represent the energy in the $\Lambda_{c}^{+}$ rest frame and the invariant mass, respectively. The calculated method of $|A|$ and $|B|$ has been applied in $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ [BESIII:2023wrw], successfully promoting the development of effective models [Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn]. In this work, $|A|$ and $|B|$ can also be derived, and their two-dimensional displays are shown in Fig. 9. We provide only the density plots and do not present numerical results due to the complex shapes of the contours cannot be well represented by a simple confidence interval.
Figure 8: The comparison about theoretical calculations and experimental measurements of decay asymmetry parameters. Gray points are theoretical calculations [Korner:1992wi, Xu:1992vc, Cheng:1991sn, Cheng:1993gf, Zenczykowski:1993hw, Zenczykowski:1993jm, Datta:1995mn, Ivanov:1997ra, Sharma:1998rd, Geng:2019xbo, Zou:2019kzq, Zhong:2022exp, Liu:2023dvg, Geng:2023pkr, Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn] and experimental measurements [CLEO:1990unw, ARGUS:1991yzs, CLEO:1995qyd, FOCUS:2005vxq, BESIII:2019odb, Belle:2022uod, Belle:2022bsi, LHCb:2024tnq, ParticleDataGroup:2024cfk] are marked in blue. Red points and yellow shadows represent the results obtained in this work. In the legend, the names of the theoretical studies and experimental measurements match the colors used in the figures, arranged in chronological order. For each theoretical study, the main theoretical models used are also specified.
Figure 9: The two-dimensional density of $|A|$ and $|B|$. Markers are theoretical calculations [Xu:1992vc, Cheng:1991sn, Cheng:1993gf, Ivanov:1997ra, Zou:2019kzq, Liu:2023dvg, Geng:2019xbo, Zhong:2022exp, Zhong:2024qqs, Cheng:2024lsn] and experimental measurements [CLEO:1995qyd], and shadows representing the distribution of $|A|$ and $|B|$ obtained in this work, $G_{F}$ is Fermi coupling constant, $c$ and $d$ means whether the model uses the results from Belle experiment with the BFs and decay asymmetry parameters of $\Xi_{c}^{0}\to\Xi^{0}\pi^{0}$, $\Xi^{0}\eta$, and $\Xi^{0}\eta^{\prime}$. The contours correspond to 68.2%, 95.4%, 99.7% confidence levels, respectively.
A secondary fit is also performed where the CPC\!P constraint is removed and is used to test the SM, and the results are shown in Appendix B. The $\alpha$-induced CPC\!P observable, weak and strong phase shifts are constructed in $A^{\alpha_{BP}}_{C\!P}=\frac{\alpha_{BP}+\bar{\alpha}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\phi_{C\!P}=\frac{\beta_{BP}+\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\Delta_{S}=\frac{\beta_{BP}-\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, where $\beta_{BP}=\sqrt{1-\alpha_{BP}^{2}}\sin\Delta_{BP}$ and $\bar{\beta}_{BP}=\sqrt{1-\bar{\alpha}_{BP}^{2}}\sin\bar{\Delta}_{BP}$. The first two observables, $A^{\alpha_{BP}}_{C\!P}$ and $\tan\phi_{C\!P}$, are CPC\!P-violating quantities, corresponding to the two constraints $\alpha_{BP}=-\bar{\alpha}_{BP}$ and $\Delta_{BP}=-\bar{\Delta}_{BP}$ imposed in the CPC\!P-conserving baseline fit, while $\tan\Delta_{S}$ is a CPC\!P-conserving observable. The $\phi_{C\!P}$ and $\Delta_{S}$ denote the weak interaction phase shift between $P$-wave and $S$-wave introduced by phase angle in the Cabibbo-Kobayashi-Maskawa (CKM) matrix and the strong phase shift. The results are summarized in Table 4. Some of the parameters in the table are derived from the fit results, and their uncertainties account for correlations among the fit parameters. The full correlation matrix is provided in Table A2 of Appendix A.
Table 4: Results of decay observables, where $\langle\alpha_{BP}\rangle$, $\langle\Delta_{BP}\rangle$, and $\delta_{p}-\delta_{s}$ are measured based on the assumption of CPC\!P conservation, while others are under the assumption of CPC\!P violation. Terms that could not be extracted are denoted by “”. All angular values in the table are expressed in radians. Among the eight parameters, $\langle\alpha_{BP}\rangle$ and $\langle\Delta_{BP}\rangle$ are directly obtained from the fit, $\langle\beta_{BP}\rangle$, $\langle\gamma_{BP}\rangle$, and $\delta_{p}-\delta_{s}$ are quantities derived from them. And $A^{\alpha_{BP}}_{C\!P}$, $\text{tan}\phi_{C\!P}$, and $\text{tan}\Delta_{S}$ are derived from the results given in Appendix B.
Parameter $\phantom{-}\Lambda_{c}^{+}\to pK_{S}^{0}$ $\phantom{-}\Lambda_{c}^{+}\to\Lambda\pi^{+}$ $\phantom{-}\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ $\phantom{-}\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$
$\langle\alpha_{BP}\rangle$ $0.918^{+0.133}_{-0.082}\pm 0.031$ $0.790\pm 0.032\pm 0.009$ $0.502\pm 0.080\pm 0.009$ $0.590\pm 0.049\pm 0.022$
$\langle\Delta_{BP}\rangle$ $\phantom{-}0.637\pm 0.444\pm 0.014$ $\phantom{-}2.190\pm 0.730\pm 0.029$ $\phantom{-}1.901\pm 0.603\pm 0.040$
$\langle\beta_{BP}\rangle$ $\phantom{-}0.365^{+0.173}_{-0.246}\pm 0.010$ $\phantom{-}0.704^{+0.143}_{-0.480}\pm 0.015$ $\phantom{-}0.764^{+0.051}_{-0.237}\pm 0.018$
$\langle\gamma_{BP}\rangle$ $\phantom{-}0.637^{+0.103}_{-0.202}\pm 0.011$ $0.502^{+0.591}_{-0.303}\pm 0.021$ $0.262^{+0.478}_{-0.383}\pm 0.031$
$\delta_{p}-\delta_{s}$ $\phantom{-}2.71^{+0.28}_{-0.17}\pm 0.02$ $\phantom{-}2.19^{+0.49}_{-0.13}\pm 0.02$ $\phantom{-}2.23^{+0.19}_{-0.06}\pm 0.03$
$A^{\alpha_{BP}}_{C\!P}$ $\phantom{-}0.079^{+0.115}_{-0.101}\pm 0.019$ $\phantom{-}0.002\pm 0.047\pm 0.017$ $\phantom{-}0.206^{+0.188}_{-0.156}\pm 0.028$ $0.086\pm 0.081\pm 0.029$
$\text{tan}\phi_{C\!P}$ $\phantom{-}0.232\pm 0.242\pm 0.025$ $\phantom{-}0.393\pm 0.651\pm 0.042$ $0.007\pm 0.474\pm 0.034$
$\text{tan}\Delta_{S}$ $0.475\pm 0.242\pm 0.029$ $1.411\pm 0.672\pm 0.062$ $1.297\pm 0.478\pm 0.068$
V Discussion
The effect of transverse polarization of $\Lambda_{c}^{+}$ baryons is discovered for the first time by analyzing $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at the CM energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The observed transverse polarization reveals that the electric and magnetic form factors of the $\Lambda^{+}_{c}$ carry a relative phase, reflecting complex dynamics in its electromagnetic structure. For the first time, we further mapped how this phase varies with collision energy, revealing clear deviations from theoretical expectations.
The vector meson dominance (VMD) model offers a framework to describe how the production cross sections of the baryon-antibaryon pairs via electron-positron annihilations change with energy [Li:2021lvs, Chen:2023oqs, Chen:2024luh], assuming that the initial electromagnatic process dominates with virtual photon coupling to the produced hadrons primarily through intermediate vector mesons. This approach has also been successfully applied to describe nucleon form factors [Yang:2019mzq]. In 2021, a VMD model incorporating the vector charmonium states was proposed [Wan:2021ncg] to describe the Born cross sections of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ measured from Belle [Belle:2008xmh] and BESIII [BESIII:2017kqg]. That time Belle data determined the overall lineshape, and the form factors $|G_{E}/G_{M}|$ in the VMD model were less constrained with sparse BESIII data. Two years later, BESIII reported the cross sections and the form factor ratio $|G_{E}/G_{M}|$ with much improved precisions at more energy points [BESIII:2023rwv], which constrains the lineshape largely. A subsequent VMD-based study including higher vector states such as $\psi(4500)$, $\psi(4660)$, $\psi(4790)$, and $\psi(4900)$ is drived to achieve an excellent description of the cross sections and $|G_{E}/G_{M}|$ at different energy points [Chen:2023oqs]. Unfortunately, neither calculation [Wan:2021ncg, Chen:2023oqs] reproduces the measured phase difference between $G_{E}$ and $G_{M}$, indicating a persistent discrepancy between theory and experiment. This suggests that the internal structure of the $\Lambda_{c}^{+}$ and its production dynamics in $e^{+}e^{-}$ annihilation are not yet fully understood.
The observed transverse polarization enables precise measurements of the $\Lambda_{c}^{+}$ decay asymmetry parameters [Li:2025nzx]. The key two-body decays $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ have been studied. The measured $\alpha_{pK^{0}_{S}}$ now changes the sign of center value and is consistent with most theoretical calculations. The decay asymmetry parameters $\alpha_{\Sigma^{+}\pi^{0}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ agree with the calculation under $SU(3)$ flavor symmetry but $\alpha_{\Sigma^{+}\pi^{0}}$ is more than $2\sigma$ different than the measurements from Belle [Belle:2022bsi]. Due to the cascading decay in $\Lambda_{c}^{+}\to\Lambda\pi^{+},~\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$, we also obtain the value of $\Delta_{BP}$, and use this to derive $\beta_{BP}$. This enables us to extract the $S$- and $P$-wave phase shift, $\delta_{p}-\delta_{s}$. The results of phase shifts in this work are consistent with recent theoretical calculations [Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn] under the convention $sign=+$ as proposed in Ref. [Wang:2024wrm]. The extra parameter signsign is introduced in the unified parameterization of $S$- and $P$-wave amplitudes in baryon decays and is important when conducting global analyses of baryon polarization, involving the strong phase shifts, among different experiments and performing future searches for baryon CPC\!P violation [Wang:2024wrm]. We have also derived the dynamic parameters $|A|$ and $|B|$ for weak interactions. Previous theoretical studies on the decay of charmed baryons have also calculated $|A|$ and $|B|$, but most of them did not account for the contribution of phase shifts to the decay asymmetry parameters, which complicates the accurate calculation of these parameters. Only a few recent works have emphasized the impact of phase shifts and predicted $|A|$ and $|B|$ [Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn]. Clearly, the reliability of their calculations is improving. The results presented here supersede those in Refs. [BESIII:2019odb, BESIII:2023rwv].
A strong phase shift can be significant in physical processes and theoretically can greatly improve the CPC\!P asymmetry prediction [He:2024unv, Geng:2023pkr, Zhong:2024qqs]. In our work, we calculate $\delta_{p}-\delta_{s}$ in three channels and demonstrate the existence of a large phase shift in charmed baryon decays, as reflected with the values of $\tan\Delta_{S}$ in Table 4. Such a large phase shift can be induced by re-scattering processes and loop effects [Geng:2023pkr, Jia:2024pyb], hence indicating that final-state interactions play an important role in QCD calculations. The decays measured in this work are all CF processes, where weak phase shifts are zero. Although the CPC\!P asymmetry phenomenon is impossible in the CF processes, it can still be transferred to other Cabibbo-suppressed processes through symmetry in physics. As discussed in Ref. [He:2024unv], the re-scattering effects can enhance CPC\!P asymmetry by an order of magnitude, which makes it promising to observe CPC\!P asymmetry effect in baryon decays. In addition, based on the separate measurements of the positive and negative processes, we also test $A^{\alpha_{BP}}_{C\!P}$, strong and weak phase shifts. They are related by $A^{\alpha_{BP}}_{C\!P}=-\tan(\delta_{p}-\delta_{s})\tan(\xi_{p}-\xi_{s})$, and the results show that $A^{\alpha_{BP}}_{C\!P}$ and weak phase shifts are consistent with zero, which aligns with the expectations of the SM.
In conclusion, the observation of the transverse polarization effect of the $\Lambda_{c}^{+}$ with the BESIII detector represents a complete description in charmed baryon production in $e^{+}e^{-}$ annihilations, contributing to our understanding of its production properties. Our analysis of decay asymmetry parameters for $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ not only updates and validates previous measurements but also reveals significant strong phase shifts, providing valuable constraints on various theoretical models. Furthermore, the systematic study of production and decay properties provides a comprehensive dataset that refines global fits in flavor physics, enhances the understanding of QCD in charm sector and the capability of prediction of the CPC\!P asymmetry in charmed baryons. These findings, combined with the precise measurement of polarization dynamics, open new avenues for investigating hadronic interactions, spin effects, and charm quark dynamics. This work establishes a robust framework for future experimental programs at Belle II, LHCb, STCF [Wang:2025yjs] and other facilities, setting the stage for deeper insights into the non-perturbative aspects of QCD and better sensitivities in searching for CPC\!P violations, as well as the exploration of potential new physics.
Data availability
The data analyzed in this study are not publicly available due to collaboration policies. Access to the underlying datasets may be granted upon reasonable request and with permission from the BESIII Collaboration by contacting besiii-publications@ihep.ac.cn.
Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII (https://cstr.cn/31109.02.BEPC) and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819, 12422504, 12365015, 125B2106; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; CAS under Contract No. YSBR-101; 100 Talents Program of CAS; Fundamental Research Funds for the Central Universities, Lanzhou University under Contracts Nos. lzujbky-2025-ytB01, lzujbky-2023-stlt01, lzujbky-2023-it32, University of Chinese Academy of Sciences; The Natural Science Foundation of Inner Mongolia Autonomous Region No. 2023QN01011; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; ERC under Contract No. 758462; German Research Foundation DFG under Contract No. FOR5327; Istituto Nazionale di Fisica Nucleare, Italy; Knut and Alice Wallenberg Foundation under Contracts Nos. 2021.0174, 2021.0299; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; Polish National Science Centre under Contract No. 2024/53/B/ST2/00975; STFC (United Kingdom); Swedish Research Council under Contract No. 2019.04595; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
Appendix A Correlation Matrix of the Fit Parameters
In this appendix, we provide the full correlation matrices shown in Table A1 under the baryon CPC\!P-conserving assumption and Table A2 without assumption of baryon CPC\!P conservation. These correlations are taken into account in the calculation of the derived quantities of $\langle\beta_{BP}\rangle$, $\langle\gamma_{BP}\rangle$, $\delta_{p}-\delta_{s}$, $A^{\alpha_{BP}}_{C\!P}$, $\tan\phi_{C\!P}$, and $\tan\Delta_{S}$.
Table A1: Correlation matrix under the assumption of baryon CPC\!P conservation. Some negligible terms are denoted by “neg”.
Parameters $\alpha_{pK^{0}_{S}}$ $\alpha_{\Lambda\pi^{+}}$ $\Delta_{\Lambda\pi^{+}}$ $\alpha_{\Sigma^{0}\pi^{+}}$ $\Delta_{\Sigma^{0}\pi^{+}}$ $\alpha_{\Sigma^{+}\pi^{0}}$ $\Delta_{\Sigma^{+}\pi^{0}}$
$\alpha_{pK^{0}_{S}}$ $1.000$ $0.024$ $0.004$ $0.011$ 0.007-0.007 0.004-0.004 0.004-0.004
$\alpha_{\Lambda\pi^{+}}$ $1.000$ 0.006-0.006 neg $0.001$ 0.001-0.001 $0.004$
$\Delta_{\Lambda\pi^{+}}$ $1.000$ $0.008$ $0.004$ neg $0.013$
$\alpha_{\Sigma^{0}\pi^{+}}$ $1.000$ 0.022-0.022 0.001-0.001 neg
$\Delta_{\Sigma^{0}\pi^{+}}$ $1.000$ neg $0.008$
$\alpha_{\Sigma^{+}\pi^{0}}$ $1.000$ $0.022$
$\Delta_{\Sigma^{+}\pi^{0}}$ $1.000$
Table A2: Correlation matrix without assumption of baryon CPC\!P conservation. baryon charge-separated case. Some negligible terms are denoted by “neg”.
Parameters $\alpha_{pK^{0}_{S}}$ $\bar{\alpha}_{pK^{0}_{S}}$ $\alpha_{\Lambda\pi^{+}}$ $\Delta_{\Lambda\pi^{+}}$ $\bar{\alpha}_{\Lambda\pi^{+}}$ $\bar{\Delta}_{\Lambda\pi^{+}}$ $\alpha_{\Sigma^{0}\pi^{+}}$ $\Delta_{\Sigma^{0}\pi^{+}}$ $\bar{\alpha}_{\Sigma^{0}\pi^{+}}$ $\bar{\Delta}_{\Sigma^{0}\pi^{+}}$ $\alpha_{\Sigma^{+}\pi^{0}}$ $\Delta_{\Sigma^{+}\pi^{0}}$ $\bar{\alpha}_{\Sigma^{+}\pi^{0}}$ $\bar{\Delta}_{\Sigma^{+}\pi^{0}}$
$\alpha_{pK^{0}_{S}}$ $1.000$ 0.141-0.141 $0.021$ $0.006$ 0.005-0.005 $0.002$ $0.008$ $0.011$ $0.008$ $0.009$ $0.001$ $0.003$ $0.003$ $0.003$
$\bar{\alpha}_{pK^{0}_{S}}$ $1.000$ 0.024-0.024 $0.001$ $0.006$ $0.021$ $0.001$ 0.008-0.008 $0.015$ 0.005-0.005 $0.001$ $0.014$ 0.009-0.009 $0.014$
$\alpha_{\Lambda\pi^{+}}$ $1.000$ 0.003-0.003 0.001-0.001 0.005-0.005 0.001-0.001 $0.003$ neg neg 0.001-0.001 $0.001$ $0.001$ 0.003-0.003
$\Delta_{\Lambda\pi^{+}}$ $1.000$ $0.002$ $0.026$ neg neg neg $0.012$ 0.001-0.001 $0.004$ 0.002-0.002 0.002-0.002
$\bar{\alpha}_{\Lambda\pi^{+}}$ $1.000$ 0.006-0.006 0.002-0.002 neg $0.001$ $0.001$ neg 0.002-0.002 neg $0.002$
$\bar{\Delta}_{\Lambda\pi^{+}}$ $1.000$ 0.001-0.001 0.007-0.007 $0.009$ $0.006$ $0.004$ $0.003$ 0.003-0.003 $0.023$
$\alpha_{\Sigma^{0}\pi^{+}}$ $1.000$ $0.001$ $0.002$ 0.004-0.004 neg $0.002$ 0.002-0.002 0.004-0.004
$\Delta_{\Sigma^{0}\pi^{+}}$ $1.000$ $0.001$ 0.001-0.001 0.002-0.002 0.002-0.002 $0.002$ 0.016-0.016
$\bar{\alpha}_{\Sigma^{0}\pi^{+}}$ $1.000$ 0.019-0.019 $0.001$ $0.004$ 0.003-0.003 $0.005$
$\bar{\Delta}_{\Sigma^{0}\pi^{+}}$ $1.000$ 0.001-0.001 0.001-0.001 neg $0.007$
$\alpha_{\Sigma^{+}\pi^{0}}$ $1.000$ $0.041$ neg $0.005$
$\Delta_{\Sigma^{+}\pi^{0}}$ $1.000$ 0.001-0.001 0.007-0.007
$\bar{\alpha}_{\Sigma^{+}\pi^{0}}$ $1.000$ 0.048-0.048
$\bar{\Delta}_{\Sigma^{+}\pi^{0}}$ $1.000$
Appendix B Results in the Baryon Charge-separated case
The fitted results of $\alpha_{BP}$, $\Delta_{BP}$, $\bar{\alpha}_{BP}$, and $\bar{\Delta}_{BP}$ are summarized in Table B1, which are measured without assumption of baryon CPC\!P conservation.
Table B1: Results of $\alpha_{BP}$, $\Delta_{BP}$, $\bar{\alpha}_{BP}$, and $\bar{\Delta}_{BP}$. The unit of $\Delta_{BP}$, and $\bar{\Delta}_{BP}$ is rad, and $\alpha_{BP}$ and $\bar{\alpha}_{BP}$ are dimensionless numbers.
Processes $\alpha_{BP}$ $\Delta_{BP}$ $\bar{\alpha}_{BP}$ $\bar{\Delta}_{BP}$
$\Lambda_{c}^{+}\to pK^{0}_{S}$ $0.994^{+0.203}_{-0.006}\pm 0.014$ $0.848^{+0.152}_{-0.179}\pm 0.031$
$\Lambda_{c}^{+}\to\Lambda\pi^{+}$ $0.796\pm 0.054\pm 0.008$ $0.325\pm 0.631\pm 0.017$ $0.793\pm 0.051\pm 0.026$ $1.171\pm 0.769\pm 0.049$
$\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ $0.607\pm 0.116\pm 0.016$ $2.440\pm 1.039\pm 0.062$ $0.400\pm 0.116\pm 0.021$ $1.704\pm 1.004\pm 0.049$
$\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$ $0.542\pm 0.062\pm 0.015$ $1.974\pm 0.777\pm 0.072$ $0.644\pm 0.076\pm 0.033$ $1.578\pm 1.334\pm 0.176$
M. Ablikim1[Uncaptioned image], M. N. Achasov4,b[Uncaptioned image], P. Adlarson77[Uncaptioned image], X. C. Ai82[Uncaptioned image], R. Aliberti36[Uncaptioned image], A. Amoroso76A,76C[Uncaptioned image], Q. An73,59,†, Y. Bai58[Uncaptioned image], O. Bakina37[Uncaptioned image], Y. Ban47,g[Uncaptioned image], H.-R. Bao65[Uncaptioned image], V. Batozskaya1,45[Uncaptioned image], K. Begzsuren33, N. Berger36[Uncaptioned image], M. Berlowski45[Uncaptioned image], M. Bertani29A[Uncaptioned image], D. Bettoni30A[Uncaptioned image], F. Bianchi76A,76C[Uncaptioned image], E. Bianco76A,76C, A. Bortone76A,76C[Uncaptioned image], I. Boyko37[Uncaptioned image], R. A. Briere5[Uncaptioned image], A. Brueggemann70[Uncaptioned image], H. Cai78[Uncaptioned image], M. H. Cai39,j,k[Uncaptioned image], X. Cai1,59[Uncaptioned image], A. Calcaterra29A[Uncaptioned image], G. F. Cao1,65[Uncaptioned image], N. Cao1,65[Uncaptioned image], S. A. Cetin63A[Uncaptioned image], X. Y. Chai47,g[Uncaptioned image], J. F. Chang1,59[Uncaptioned image], G. R. Che44[Uncaptioned image], Y. Z. Che1,59,65[Uncaptioned image], C. H. Chen9[Uncaptioned image], Chao Chen56[Uncaptioned image], G. Chen1[Uncaptioned image], H. S. Chen1,65[Uncaptioned image], H. Y. Chen21[Uncaptioned image], M. L. Chen1,59,65[Uncaptioned image], S. J. Chen43[Uncaptioned image], S. L. Chen46[Uncaptioned image], S. M. Chen62[Uncaptioned image], T. Chen1,65[Uncaptioned image], X. R. Chen32,65[Uncaptioned image], X. T. Chen1,65[Uncaptioned image], X. Y. Chen12,f[Uncaptioned image], Y. B. Chen1,59[Uncaptioned image], Y. Q. Chen35[Uncaptioned image], Y. Q. Chen16[Uncaptioned image], Z. Chen25[Uncaptioned image], Z. J. Chen26,h[Uncaptioned image], Z. K. Chen60[Uncaptioned image], S. K. Choi10[Uncaptioned image], X. Chu12,f[Uncaptioned image], G. Cibinetto30A[Uncaptioned image], F. Cossio76C[Uncaptioned image], J. Cottee-Meldrum64[Uncaptioned image], J. J. Cui51[Uncaptioned image], H. L. Dai1,59[Uncaptioned image], J. P. Dai80[Uncaptioned image], A. Dbeyssi19, R. E. de Boer3[Uncaptioned image], D. Dedovich37[Uncaptioned image], C. Q. Deng74[Uncaptioned image], Z. Y. Deng1[Uncaptioned image], A. Denig36[Uncaptioned image], I. Denysenko37[Uncaptioned image], M. Destefanis76A,76C[Uncaptioned image], F. De Mori76A,76C[Uncaptioned image], B. Ding68,1[Uncaptioned image], X. X. Ding47,g[Uncaptioned image], Y. Ding41[Uncaptioned image], Y. Ding35[Uncaptioned image], Y. X. Ding31[Uncaptioned image], J. Dong1,59[Uncaptioned image], L. Y. Dong1,65[Uncaptioned image], M. Y. Dong1,59,65[Uncaptioned image], X. Dong78[Uncaptioned image], M. C. Du1[Uncaptioned image], S. X. Du82[Uncaptioned image], S. X. Du12,f[Uncaptioned image], Y. Y. Duan56[Uncaptioned image], P. Egorov37,a[Uncaptioned image], G. F. Fan43[Uncaptioned image], J. J. Fan20[Uncaptioned image], Y. H. Fan46[Uncaptioned image], J. Fang1,59[Uncaptioned image], J. Fang60[Uncaptioned image], S. S. Fang1,65[Uncaptioned image], W. X. Fang1[Uncaptioned image], Y. Q. Fang1,59[Uncaptioned image], R. Farinelli30A[Uncaptioned image], L. Fava76B,76C[Uncaptioned image], F. Feldbauer3[Uncaptioned image], G. Felici29A[Uncaptioned image], C. Q. Feng73,59[Uncaptioned image], J. H. Feng16[Uncaptioned image], L. Feng39,j,k[Uncaptioned image], Q. X. Feng39,j,k[Uncaptioned image], Y. T. Feng73,59[Uncaptioned image], M. Fritsch3[Uncaptioned image], C. D. Fu1[Uncaptioned image], J. L. Fu65[Uncaptioned image], Y. W. Fu1,65[Uncaptioned image], H. Gao65[Uncaptioned image], X. B. Gao42[Uncaptioned image], Y. Gao73,59[Uncaptioned image], Y. N. Gao47,g[Uncaptioned image], Y. N. Gao20[Uncaptioned image], Y. Y. Gao31[Uncaptioned image], S. Garbolino76C[Uncaptioned image], I. Garzia30A,30B[Uncaptioned image], L. Ge58[Uncaptioned image], P. T. Ge20[Uncaptioned image], Z. W. Ge43[Uncaptioned image], C. Geng60[Uncaptioned image], E. M. Gersabeck69[Uncaptioned image], A. Gilman71[Uncaptioned image], K. Goetzen13[Uncaptioned image], J. D. Gong35[Uncaptioned image], L. Gong41[Uncaptioned image], W. X. Gong1,59[Uncaptioned image], W. Gradl36[Uncaptioned image], S. Gramigna30A,30B[Uncaptioned image], M. Greco76A,76C[Uncaptioned image], M. H. Gu1,59[Uncaptioned image], Y. T. Gu15[Uncaptioned image], C. Y. Guan1,65[Uncaptioned image], A. Q. Guo32[Uncaptioned image], L. B. Guo42[Uncaptioned image], M. J. Guo51[Uncaptioned image], R. P. Guo50[Uncaptioned image], Y. P. Guo12,f[Uncaptioned image], A. Guskov37,a[Uncaptioned image], J. Gutierrez28[Uncaptioned image], K. L. Han65[Uncaptioned image], T. T. Han1[Uncaptioned image], F. Hanisch3[Uncaptioned image], K. D. Hao73,59[Uncaptioned image], X. Q. Hao20[Uncaptioned image], F. A. Harris67[Uncaptioned image], K. K. He56[Uncaptioned image], K. L. He1,65[Uncaptioned image], F. H. Heinsius3[Uncaptioned image], C. H. Heinz36[Uncaptioned image], Y. K. Heng1,59,65[Uncaptioned image], C. Herold61[Uncaptioned image], P. C. Hong35[Uncaptioned image], G. Y. Hou1,65[Uncaptioned image], X. T. Hou1,65[Uncaptioned image], Y. R. Hou65[Uncaptioned image], Z. L. Hou1[Uncaptioned image], H. M. Hu1,65[Uncaptioned image], J. F. Hu57,i[Uncaptioned image], Q. P. Hu73,59[Uncaptioned image], S. L. Hu12,f[Uncaptioned image], T. Hu1,59,65[Uncaptioned image], Y. Hu1[Uncaptioned image], Z. M. Hu60[Uncaptioned image], G. S. Huang73,59[Uncaptioned image], K. X. Huang60[Uncaptioned image], L. Q. Huang32,65[Uncaptioned image], P. Huang43[Uncaptioned image], X. T. Huang51[Uncaptioned image], Y. P. Huang1[Uncaptioned image], Y. S. Huang60[Uncaptioned image], T. Hussain75[Uncaptioned image], N. Hüsken36[Uncaptioned image], N. in der Wiesche70[Uncaptioned image], J. Jackson28[Uncaptioned image], Q. Ji1[Uncaptioned image], Q. P. Ji20[Uncaptioned image], W. Ji1,65[Uncaptioned image], X. B. Ji1,65[Uncaptioned image], X. L. Ji1,59[Uncaptioned image], Y. Y. Ji51[Uncaptioned image], Z. K. Jia73,59[Uncaptioned image], D. Jiang1,65[Uncaptioned image], H. B. Jiang78[Uncaptioned image], P. C. Jiang47,g[Uncaptioned image], S. J. Jiang9[Uncaptioned image], T. J. Jiang17[Uncaptioned image], X. S. Jiang1,59,65[Uncaptioned image], Y. Jiang65[Uncaptioned image], J. B. Jiao51[Uncaptioned image], J. K. Jiao35[Uncaptioned image], Z. Jiao24[Uncaptioned image], S. Jin43[Uncaptioned image], Y. Jin68[Uncaptioned image], M. Q. Jing1,65[Uncaptioned image], X. M. Jing65[Uncaptioned image], T. Johansson77[Uncaptioned image], S. Kabana34[Uncaptioned image], N. Kalantar-Nayestanaki66[Uncaptioned image], X. L. Kang9[Uncaptioned image], X. S. Kang41[Uncaptioned image], M. Kavatsyuk66[Uncaptioned image], B. C. Ke82[Uncaptioned image], V. Khachatryan28[Uncaptioned image], A. Khoukaz70[Uncaptioned image], R. Kiuchi1, O. B. Kolcu63A[Uncaptioned image], B. Kopf3[Uncaptioned image], M. Kuessner3[Uncaptioned image], X. Kui1,65[Uncaptioned image], N. Kumar27[Uncaptioned image], A. Kupsc45,77[Uncaptioned image], W. Kühn38[Uncaptioned image], Q. Lan74[Uncaptioned image], W. N. Lan20[Uncaptioned image], T. T. Lei73,59[Uncaptioned image], M. Lellmann36[Uncaptioned image], T. Lenz36[Uncaptioned image], C. Li48[Uncaptioned image], C. Li44[Uncaptioned image], C. H. Li40[Uncaptioned image], C. K. Li21[Uncaptioned image], D. M. Li82[Uncaptioned image], F. Li1,59[Uncaptioned image], G. Li1[Uncaptioned image], H. B. Li1,65[Uncaptioned image], H. J. Li20[Uncaptioned image], H. N. Li57,i[Uncaptioned image], Hui Li44[Uncaptioned image], J. R. Li62[Uncaptioned image], J. S. Li60[Uncaptioned image], K. Li1[Uncaptioned image], K. L. Li20[Uncaptioned image], K. L. Li39,j,k[Uncaptioned image], L. J. Li1,65[Uncaptioned image], Lei Li49[Uncaptioned image], M. H. Li44[Uncaptioned image], M. R. Li1,65[Uncaptioned image], P. L. Li65[Uncaptioned image], P. R. Li39,j,k[Uncaptioned image], Q. M. Li1,65[Uncaptioned image], Q. X. Li51[Uncaptioned image], R. Li18,32[Uncaptioned image], S. X. Li12[Uncaptioned image], T. Li51[Uncaptioned image], T. Y. Li44[Uncaptioned image], W. D. Li1,65[Uncaptioned image], W. G. Li1,†[Uncaptioned image], X. Li1,65[Uncaptioned image], X. H. Li73,59[Uncaptioned image], X. L. Li51[Uncaptioned image], X. Y. Li1,8[Uncaptioned image], X. Z. Li60[Uncaptioned image], Y. Li20[Uncaptioned image], Y. G. Li47,g[Uncaptioned image], Y. P. Li35[Uncaptioned image], Z. J. Li60[Uncaptioned image], Z. Y. Li80[Uncaptioned image], H. Liang73,59[Uncaptioned image], Y. F. Liang55[Uncaptioned image], Y. T. Liang32,65[Uncaptioned image], G. R. Liao14[Uncaptioned image], L. B. Liao60[Uncaptioned image], M. H. Liao60[Uncaptioned image], Y. P. Liao1,65[Uncaptioned image], J. Libby27[Uncaptioned image], A. Limphirat61[Uncaptioned image], C. C. Lin56[Uncaptioned image], D. X. Lin32,65[Uncaptioned image], L. Q. Lin40[Uncaptioned image], T. Lin1[Uncaptioned image], B. J. Liu1[Uncaptioned image], B. X. Liu78[Uncaptioned image], C. Liu35[Uncaptioned image], C. X. Liu1[Uncaptioned image], F. Liu1[Uncaptioned image], F. H. Liu54[Uncaptioned image], Feng Liu6[Uncaptioned image], G. M. Liu57,i[Uncaptioned image], H. Liu39,j,k[Uncaptioned image], H. B. Liu15[Uncaptioned image], H. H. Liu1[Uncaptioned image], H. M. Liu1,65[Uncaptioned image], Huihui Liu22[Uncaptioned image], J. B. Liu73,59[Uncaptioned image], J. J. Liu21[Uncaptioned image], K. Liu39,j,k[Uncaptioned image], K. Liu74[Uncaptioned image], K. Y. Liu41[Uncaptioned image], Ke Liu23[Uncaptioned image], L. C. Liu44[Uncaptioned image], Lu Liu44[Uncaptioned image], M. H. Liu12,f[Uncaptioned image], P. L. Liu1[Uncaptioned image], Q. Liu65[Uncaptioned image], S. B. Liu73,59[Uncaptioned image], T. Liu12,f[Uncaptioned image], W. K. Liu44[Uncaptioned image], W. M. Liu73,59[Uncaptioned image], W. T. Liu40[Uncaptioned image], X. Liu39,j,k[Uncaptioned image], X. Liu40[Uncaptioned image], X. K. Liu39,j,k[Uncaptioned image], X. L. Liu12,f[Uncaptioned image], X. Y. Liu78[Uncaptioned image], Y. Liu39,j,k[Uncaptioned image], Y. Liu82[Uncaptioned image], Yuan Liu82[Uncaptioned image], Y. B. Liu44[Uncaptioned image], Z. A. Liu1,59,65[Uncaptioned image], Z. D. Liu9[Uncaptioned image], Z. Q. Liu51[Uncaptioned image], X. C. Lou1,59,65[Uncaptioned image], F. X. Lu60[Uncaptioned image], H. J. Lu24[Uncaptioned image], J. G. Lu1,59[Uncaptioned image], X. L. Lu16[Uncaptioned image], Y. Lu7[Uncaptioned image], Y. H. Lu1,65[Uncaptioned image], Y. P. Lu1,59[Uncaptioned image], Z. H. Lu1,65[Uncaptioned image], C. L. Luo42[Uncaptioned image], J. R. Luo60[Uncaptioned image], J. S. Luo1,65[Uncaptioned image], M. X. Luo81, T. Luo12,f[Uncaptioned image], X. L. Luo1,59[Uncaptioned image], Z. Y. Lv23[Uncaptioned image], X. R. Lyu65,o[Uncaptioned image], Y. F. Lyu44[Uncaptioned image], Y. H. Lyu82[Uncaptioned image], F. C. Ma41[Uncaptioned image], H. L. Ma1[Uncaptioned image], J. L. Ma1,65[Uncaptioned image], L. L. Ma51[Uncaptioned image], L. R. Ma68[Uncaptioned image], Q. M. Ma1[Uncaptioned image], R. Q. Ma1,65[Uncaptioned image], R. Y. Ma20[Uncaptioned image], T. Ma73,59[Uncaptioned image], X. T. Ma1,65[Uncaptioned image], X. Y. Ma1,59[Uncaptioned image], Y. M. Ma32[Uncaptioned image], F. E. Maas19[Uncaptioned image], I. MacKay71[Uncaptioned image], M. Maggiora76A,76C[Uncaptioned image], S. Malde71[Uncaptioned image], Q. A. Malik75[Uncaptioned image], H. X. Mao39,j,k[Uncaptioned image], Y. J. Mao47,g[Uncaptioned image], Z. P. Mao1[Uncaptioned image], S. Marcello76A,76C[Uncaptioned image], A. Marshall64[Uncaptioned image], F. M. Melendi30A,30B[Uncaptioned image], Y. H. Meng65[Uncaptioned image], Z. X. Meng68[Uncaptioned image], G. Mezzadri30A[Uncaptioned image], H. Miao1,65[Uncaptioned image], T. J. Min43[Uncaptioned image], R. E. Mitchell28[Uncaptioned image], X. H. Mo1,59,65[Uncaptioned image], B. Moses28[Uncaptioned image], N. Yu. Muchnoi4,b[Uncaptioned image], J. Muskalla36[Uncaptioned image], Y. Nefedov37[Uncaptioned image], F. Nerling19,d[Uncaptioned image], L. S. Nie21[Uncaptioned image], I. B. Nikolaev4,b, Z. Ning1,59[Uncaptioned image], S. Nisar11,l, Q. L. Niu39,j,k[Uncaptioned image], W. D. Niu12,f[Uncaptioned image], C. Normand64[Uncaptioned image], S. L. Olsen10,65[Uncaptioned image], Q. Ouyang1,59,65[Uncaptioned image], S. Pacetti29B,29C[Uncaptioned image], X. Pan56[Uncaptioned image], Y. Pan58[Uncaptioned image], A. Pathak10[Uncaptioned image], Y. P. Pei73,59[Uncaptioned image], M. Pelizaeus3[Uncaptioned image], H. P. Peng73,59[Uncaptioned image], X. J. Peng39,j,k[Uncaptioned image], Y. Y. Peng39,j,k[Uncaptioned image], K. Peters13,d[Uncaptioned image], K. Petridis64[Uncaptioned image], J. L. Ping42[Uncaptioned image], R. G. Ping1,65[Uncaptioned image], S. Plura36[Uncaptioned image], V. Prasad35[Uncaptioned image], F. Z. Qi1[Uncaptioned image], H. R. Qi62[Uncaptioned image], M. Qi43[Uncaptioned image], S. Qian1,59[Uncaptioned image], W. B. Qian65[Uncaptioned image], C. F. Qiao65[Uncaptioned image], J. H. Qiao20[Uncaptioned image], J. J. Qin74[Uncaptioned image], J. L. Qin56[Uncaptioned image], L. Q. Qin14[Uncaptioned image], L. Y. Qin73,59[Uncaptioned image], P. B. Qin74[Uncaptioned image], X. P. Qin12,f[Uncaptioned image], X. S. Qin51[Uncaptioned image], Z. H. Qin1,59[Uncaptioned image], J. F. Qiu1[Uncaptioned image], Z. H. Qu74[Uncaptioned image], J. Rademacker64[Uncaptioned image], C. F. Redmer36[Uncaptioned image], A. Rivetti76C[Uncaptioned image], M. Rolo76C[Uncaptioned image], G. Rong1,65[Uncaptioned image], S. S. Rong1,65[Uncaptioned image], F. Rosini29B,29C[Uncaptioned image], Ch. Rosner19[Uncaptioned image], M. Q. Ruan1,59[Uncaptioned image], N. Salone45[Uncaptioned image], A. Sarantsev37,c[Uncaptioned image], Y. Schelhaas36[Uncaptioned image], K. Schoenning77[Uncaptioned image], M. Scodeggio30A[Uncaptioned image], K. Y. Shan12,f[Uncaptioned image], W. Shan25[Uncaptioned image], X. Y. Shan73,59[Uncaptioned image], Z. J. Shang39,j,k[Uncaptioned image], J. F. Shangguan17[Uncaptioned image], L. G. Shao1,65[Uncaptioned image], M. Shao73,59[Uncaptioned image], C. P. Shen12,f[Uncaptioned image], H. F. Shen1,8[Uncaptioned image], W. H. Shen65[Uncaptioned image], X. Y. Shen1,65[Uncaptioned image], B. A. Shi65[Uncaptioned image], H. Shi73,59[Uncaptioned image], J. L. Shi12,f[Uncaptioned image], J. Y. Shi1[Uncaptioned image], S. Y. Shi74[Uncaptioned image], X. Shi1,59[Uncaptioned image], H. L. Song73,59[Uncaptioned image], J. J. Song20[Uncaptioned image], T. Z. Song60[Uncaptioned image], W. M. Song35[Uncaptioned image], Y. J. Song12,f[Uncaptioned image], Y. X. Song47,g,m[Uncaptioned image], S. Sosio76A,76C[Uncaptioned image], S. Spataro76A,76C[Uncaptioned image], F. Stieler36[Uncaptioned image], S. S Su41[Uncaptioned image], Y. J. Su65[Uncaptioned image], G. B. Sun78[Uncaptioned image], G. X. Sun1[Uncaptioned image], H. Sun65[Uncaptioned image], H. K. Sun1[Uncaptioned image], J. F. Sun20[Uncaptioned image], K. Sun62[Uncaptioned image], L. Sun78[Uncaptioned image], S. S. Sun1,65[Uncaptioned image], T. Sun52,e[Uncaptioned image], Y. C. Sun78[Uncaptioned image], Y. H. Sun31[Uncaptioned image], Y. J. Sun73,59[Uncaptioned image], Y. Z. Sun1[Uncaptioned image], Z. Q. Sun1,65[Uncaptioned image], Z. T. Sun51[Uncaptioned image], C. J. Tang55, G. Y. Tang1[Uncaptioned image], J. Tang60[Uncaptioned image], J. J. Tang73,59[Uncaptioned image], L. F. Tang40[Uncaptioned image], Y. A. Tang78[Uncaptioned image], L. Y. Tao74[Uncaptioned image], M. Tat71[Uncaptioned image], J. X. Teng73,59[Uncaptioned image], J. Y. Tian73,59[Uncaptioned image], W. H. Tian60[Uncaptioned image], Y. Tian32[Uncaptioned image], Z. F. Tian78[Uncaptioned image], I. Uman63B[Uncaptioned image], B. Wang1[Uncaptioned image], B. Wang60[Uncaptioned image], Bo Wang73,59[Uncaptioned image], C. Wang39,j,k[Uncaptioned image], C. Wang20[Uncaptioned image], Cong Wang23[Uncaptioned image], D. Y. Wang47,g[Uncaptioned image], H. J. Wang39,j,k[Uncaptioned image], J. J. Wang78[Uncaptioned image], K. Wang1,59[Uncaptioned image], L. L. Wang1[Uncaptioned image], L. W. Wang35[Uncaptioned image], M. Wang51[Uncaptioned image], M. Wang73,59[Uncaptioned image], N. Y. Wang65[Uncaptioned image], S. Wang12,f[Uncaptioned image], T. Wang12,f[Uncaptioned image], T. J. Wang44[Uncaptioned image], W. Wang60[Uncaptioned image], Wei Wang74[Uncaptioned image], W. P. Wang36,73,59,n[Uncaptioned image], X. Wang47,g[Uncaptioned image], X. F. Wang39,j,k[Uncaptioned image], X. J. Wang40[Uncaptioned image], X. L. Wang12,f[Uncaptioned image], X. N. Wang1,65[Uncaptioned image], Y. Wang62[Uncaptioned image], Y. D. Wang46[Uncaptioned image], Y. F. Wang1,8,65[Uncaptioned image], Y. H. Wang39,j,k[Uncaptioned image], Y. J. Wang73,59[Uncaptioned image], Y. L. Wang20[Uncaptioned image], Y. N. Wang78[Uncaptioned image], Y. Q. Wang1[Uncaptioned image], Yaqian Wang18[Uncaptioned image], Yi Wang62[Uncaptioned image], Yuan Wang18,32[Uncaptioned image], Z. Wang1,59[Uncaptioned image], Z. L. Wang74[Uncaptioned image], Z. L. Wang2[Uncaptioned image], Z. Q. Wang12,f[Uncaptioned image], Z. Y. Wang1,65[Uncaptioned image], D. H. Wei14[Uncaptioned image], H. R. Wei44[Uncaptioned image], F. Weidner70[Uncaptioned image], S. P. Wen1[Uncaptioned image], Y. R. Wen40[Uncaptioned image], U. Wiedner3[Uncaptioned image], G. Wilkinson71[Uncaptioned image], M. Wolke77, C. Wu40[Uncaptioned image], J. F. Wu1,8[Uncaptioned image], L. H. Wu1[Uncaptioned image], L. J. Wu1,65[Uncaptioned image], L. J. Wu20[Uncaptioned image], Lianjie Wu20[Uncaptioned image], S. G. Wu1,65[Uncaptioned image], S. M. Wu65[Uncaptioned image], X. Wu12,f[Uncaptioned image], X. H. Wu35[Uncaptioned image], Y. J. Wu32[Uncaptioned image], Z. Wu1,59[Uncaptioned image], L. Xia73,59[Uncaptioned image], X. M. Xian40[Uncaptioned image], B. H. Xiang1,65[Uncaptioned image], D. Xiao39,j,k[Uncaptioned image], G. Y. Xiao43[Uncaptioned image], H. Xiao74[Uncaptioned image], Y. L. Xiao12,f[Uncaptioned image], Z. J. Xiao42[Uncaptioned image], C. Xie43[Uncaptioned image], K. J. Xie1,65[Uncaptioned image], X. H. Xie47,g[Uncaptioned image], Y. Xie51[Uncaptioned image], Y. G. Xie1,59[Uncaptioned image], Y. H. Xie6[Uncaptioned image], Z. P. Xie73,59[Uncaptioned image], T. Y. Xing1,65[Uncaptioned image], C. F. Xu1,65, C. J. Xu60[Uncaptioned image], G. F. Xu1[Uncaptioned image], H. Y. Xu68,2[Uncaptioned image], H. Y. Xu2[Uncaptioned image], M. Xu73,59[Uncaptioned image], Q. J. Xu17[Uncaptioned image], Q. N. Xu31[Uncaptioned image], T. D. Xu74[Uncaptioned image], W. Xu1[Uncaptioned image], W. L. Xu68[Uncaptioned image], X. P. Xu56[Uncaptioned image], Y. Xu41[Uncaptioned image], Y. Xu12,f[Uncaptioned image], Y. C. Xu79[Uncaptioned image], Z. S. Xu65[Uncaptioned image], F. Yan12,f[Uncaptioned image], H. Y. Yan40[Uncaptioned image], L. Yan12,f[Uncaptioned image], W. B. Yan73,59[Uncaptioned image], W. C. Yan82[Uncaptioned image], W. H. Yan6[Uncaptioned image], W. P. Yan20[Uncaptioned image], X. Q. Yan1,65[Uncaptioned image], H. J. Yang52,e[Uncaptioned image], H. L. Yang35[Uncaptioned image], H. X. Yang1[Uncaptioned image], J. H. Yang43[Uncaptioned image], R. J. Yang20[Uncaptioned image], T. Yang1[Uncaptioned image], Y. Yang12,f[Uncaptioned image], Y. F. Yang44[Uncaptioned image], Y. H. Yang43[Uncaptioned image], Y. Q. Yang9[Uncaptioned image], Y. X. Yang1,65[Uncaptioned image], Y. Z. Yang20[Uncaptioned image], M. Ye1,59[Uncaptioned image], M. H. Ye8,†[Uncaptioned image], Z. J. Ye57,i[Uncaptioned image], Junhao Yin44[Uncaptioned image], Z. Y. You60[Uncaptioned image], B. X. Yu1,59,65[Uncaptioned image], C. X. Yu44[Uncaptioned image], G. Yu13[Uncaptioned image], J. S. Yu26,h[Uncaptioned image], L. Q. Yu12,f[Uncaptioned image], M. C. Yu41[Uncaptioned image], T. Yu74[Uncaptioned image], X. D. Yu47,g[Uncaptioned image], Y. C. Yu82[Uncaptioned image], C. Z. Yuan1,65[Uncaptioned image], H. Yuan1,65[Uncaptioned image], J. Yuan35[Uncaptioned image], J. Yuan46[Uncaptioned image], L. Yuan2[Uncaptioned image], S. C. Yuan1,65[Uncaptioned image], X. Q. Yuan1[Uncaptioned image], Y. Yuan1,65[Uncaptioned image], Z. Y. Yuan60[Uncaptioned image], C. X. Yue40[Uncaptioned image], Ying Yue20[Uncaptioned image], A. A. Zafar75[Uncaptioned image], S. H. Zeng64[Uncaptioned image], X. Zeng12,f[Uncaptioned image], Y. Zeng26,h, Yujie Zeng60[Uncaptioned image], Y. J. Zeng1,65[Uncaptioned image], X. Y. Zhai35[Uncaptioned image], Y. H. Zhan60[Uncaptioned image], Shunan Zhang71[Uncaptioned image], A. Q. Zhang1,65[Uncaptioned image], B. L. Zhang1,65[Uncaptioned image], B. X. Zhang1[Uncaptioned image], D. H. Zhang44[Uncaptioned image], G. Y. Zhang20[Uncaptioned image], G. Y. Zhang1,65[Uncaptioned image], H. Zhang73,59[Uncaptioned image], H. Zhang82[Uncaptioned image], H. C. Zhang1,59,65[Uncaptioned image], H. H. Zhang60[Uncaptioned image], H. Q. Zhang1,59,65[Uncaptioned image], H. R. Zhang73,59[Uncaptioned image], H. Y. Zhang1,59[Uncaptioned image], Jin Zhang82[Uncaptioned image], J. Zhang60[Uncaptioned image], J. J. Zhang53[Uncaptioned image], J. L. Zhang21[Uncaptioned image], J. Q. Zhang42[Uncaptioned image], J. S. Zhang12,f[Uncaptioned image], J. W. Zhang1,59,65[Uncaptioned image], J. X. Zhang39,j,k[Uncaptioned image], J. Y. Zhang1[Uncaptioned image], J. Z. Zhang1,65[Uncaptioned image], Jianyu Zhang65[Uncaptioned image], L. M. Zhang62[Uncaptioned image], Lei Zhang43[Uncaptioned image], N. Zhang82[Uncaptioned image], P. Zhang1,8[Uncaptioned image], Q. Zhang20[Uncaptioned image], Q. Y. Zhang35[Uncaptioned image], R. Y. Zhang39,j,k[Uncaptioned image], S. H. Zhang1,65[Uncaptioned image], Shulei Zhang26,h[Uncaptioned image], X. M. Zhang1[Uncaptioned image], X. Y Zhang41[Uncaptioned image], X. Y. Zhang51[Uncaptioned image], Y. Zhang1[Uncaptioned image], Y. Zhang74[Uncaptioned image], Y. T. Zhang82[Uncaptioned image], Y. H. Zhang1,59[Uncaptioned image], Y. M. Zhang40[Uncaptioned image], Y. P. Zhang73,59[Uncaptioned image], Z. D. Zhang1[Uncaptioned image], Z. H. Zhang1[Uncaptioned image], Z. L. Zhang35[Uncaptioned image], Z. L. Zhang56[Uncaptioned image], Z. X. Zhang20[Uncaptioned image], Z. Y. Zhang78[Uncaptioned image], Z. Y. Zhang44[Uncaptioned image], Z. Z. Zhang46[Uncaptioned image], Zh. Zh. Zhang20[Uncaptioned image], G. Zhao1[Uncaptioned image], J. Y. Zhao1,65[Uncaptioned image], J. Z. Zhao1,59[Uncaptioned image], L. Zhao1[Uncaptioned image], L. Zhao73,59[Uncaptioned image], M. G. Zhao44[Uncaptioned image], N. Zhao80[Uncaptioned image], R. P. Zhao65[Uncaptioned image], S. J. Zhao82[Uncaptioned image], Y. B. Zhao1,59[Uncaptioned image], Y. L. Zhao56[Uncaptioned image], Y. X. Zhao32,65[Uncaptioned image], Z. G. Zhao73,59[Uncaptioned image], A. Zhemchugov37,a[Uncaptioned image], B. Zheng74[Uncaptioned image], B. M. Zheng35[Uncaptioned image], J. P. Zheng1,59[Uncaptioned image], W. J. Zheng1,65[Uncaptioned image], X. R. Zheng20[Uncaptioned image], Y. H. Zheng65,o[Uncaptioned image], B. Zhong42[Uncaptioned image], C. Zhong20[Uncaptioned image], H. Zhou36,51,n[Uncaptioned image], J. Q. Zhou35[Uncaptioned image], J. Y. Zhou35[Uncaptioned image], S. Zhou6[Uncaptioned image], X. Zhou78[Uncaptioned image], X. K. Zhou6[Uncaptioned image], X. R. Zhou73,59[Uncaptioned image], X. Y. Zhou40[Uncaptioned image], Y. X. Zhou79[Uncaptioned image], Y. Z. Zhou12,f[Uncaptioned image], A. N. Zhu65[Uncaptioned image], J. Zhu44[Uncaptioned image], K. Zhu1[Uncaptioned image], K. J. Zhu1,59,65[Uncaptioned image], K. S. Zhu12,f[Uncaptioned image], L. Zhu35[Uncaptioned image], L. X. Zhu65[Uncaptioned image], S. H. Zhu72[Uncaptioned image], T. J. Zhu12,f[Uncaptioned image], W. D. Zhu42[Uncaptioned image], W. D. Zhu12,f[Uncaptioned image], W. J. Zhu1[Uncaptioned image], W. Z. Zhu20[Uncaptioned image], Y. C. Zhu73,59[Uncaptioned image], Z. A. Zhu1,65[Uncaptioned image], X. Y. Zhuang44[Uncaptioned image], J. H. Zou1[Uncaptioned image], J. Zu73,59[Uncaptioned image]
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Guangxi University of Science and Technology, Liuzhou 545006, People’s Republic of China
17 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
18 Hebei University, Baoding 071002, People’s Republic of China
19 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
20 Henan Normal University, Xinxiang 453007, People’s Republic of China
21 Henan University, Kaifeng 475004, People’s Republic of China
22 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
23 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
24 Huangshan College, Huangshan 245000, People’s Republic of China
25 Hunan Normal University, Changsha 410081, People’s Republic of China
26 Hunan University, Changsha 410082, People’s Republic of China
27 Indian Institute of Technology Madras, Chennai 600036, India
28 Indiana University, Bloomington, Indiana 47405, USA
29 INFN Laboratori Nazionali di Frascati, (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
30 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
31 Inner Mongolia University, Hohhot 010021, People’s Republic of China
32 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
33 Institute of Physics and Technology, Mongolian Academy of Sciences, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
34 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
35 Jilin University, Changchun 130012, People’s Republic of China
36 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
37 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
38 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
39 Lanzhou University, Lanzhou 730000, People’s Republic of China
40 Liaoning Normal University, Dalian 116029, People’s Republic of China
41 Liaoning University, Shenyang 110036, People’s Republic of China
42 Nanjing Normal University, Nanjing 210023, People’s Republic of China
43 Nanjing University, Nanjing 210093, People’s Republic of China
44 Nankai University, Tianjin 300071, People’s Republic of China
45 National Centre for Nuclear Research, Warsaw 02-093, Poland
46 North China Electric Power University, Beijing 102206, People’s Republic of China
47 Peking University, Beijing 100871, People’s Republic of China
48 Qufu Normal University, Qufu 273165, People’s Republic of China
49 Renmin University of China, Beijing 100872, People’s Republic of China
50 Shandong Normal University, Jinan 250014, People’s Republic of China
51 Shandong University, Jinan 250100, People’s Republic of China
52 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
53 Shanxi Normal University, Linfen 041004, People’s Republic of China
54 Shanxi University, Taiyuan 030006, People’s Republic of China
55 Sichuan University, Chengdu 610064, People’s Republic of China
56 Soochow University, Suzhou 215006, People’s Republic of China
57 South China Normal University, Guangzhou 510006, People’s Republic of China
58 Southeast University, Nanjing 211100, People’s Republic of China
59 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
60 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
61 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
62 Tsinghua University, Beijing 100084, People’s Republic of China
63 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
64 University of Bristol, H H Wills Physics Laboratory, Tyndall Avenue, Bristol, BS8 1TL, UK
65 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
66 University of Groningen, NL-9747 AA Groningen, The Netherlands
67 University of Hawaii, Honolulu, Hawaii 96822, USA
68 University of Jinan, Jinan 250022, People’s Republic of China
69 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
70 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
71 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
72 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
73 University of Science and Technology of China, Hefei 230026, People’s Republic of China
74 University of South China, Hengyang 421001, People’s Republic of China
75 University of the Punjab, Lahore-54590, Pakistan
76 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
77 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
78 Wuhan University, Wuhan 430072, People’s Republic of China
79 Yantai University, Yantai 264005, People’s Republic of China
80 Yunnan University, Kunming 650500, People’s Republic of China
81 Zhejiang University, Hangzhou 310027, People’s Republic of China
82 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
Deceased
a Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
b Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
c Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
d Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
e Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
f Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
g Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
h Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
i Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
j Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
k Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
m Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
n Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
o Also at Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences, Hangzhou 310024, China